### v1

# Lipschitz-Guided Design of Interpolation Schedules in Generative Models

Abstract.
We study the design of interpolation schedules in the stochastic interpolants framework for flow and diffusion-based generative models. We show that while all scalar interpolation schedules achieve identical statistical efficiency under Kullback-Leibler divergence in path space after optimal diffusion coefficient tuning, their numerical efficiency can differ substantially. This observation motivates focusing on numerical properties of the resulting drift fields rather than statistical criteria for schedule design. We propose averaged squared Lipschitzness minimization as a principled criterion for numerical optimization, providing an alternative to kinetic energy minimization used in optimal transport approaches. A transfer formula is derived that enables conversion between different schedules at inference time without retraining neural networks. For Gaussian distributions, our optimized schedules achieve exponential improvements in Lipschitz constants over standard linear schedules, while for Gaussian mixtures, they reduce mode collapse in few-step sampling. We also validate our approach on high-dimensional invariant distributions from stochastic Allen-Cahn equations and Navier-Stokes equations, demonstrating robust performance improvements across resolutions.
1. Introduction
1.1. Context
Dynamics between probability measures, particularly flows and diffusion processes described by ordinary and stochastic differential equations (ODEs and SDEs), form the foundation of state-of-the-art generative modeling techniques [38, 19, 42]. These methods generate samples through an iterative refinement process that progressively eliminates noise or corruption at different scales [40, 23].
In this paper, we study the impact and design of interpolation schedules on the performance of flow and diffusion-based generative models. We work within the stochastic interpolant framework [1, 2], which provides a systematic approach for modeling noising processes through sample interpolation and enables principled construction of the corresponding generative processes. This framework connects to related concurrent work on flow matching [27] and rectified flows [28], and encompasses diffusion and score-based generative models [38, 42, 19, 41] as specific instances.
1.2. Basics of stochastic interpolants
Let $x_{1}\sim\mu^{*}$, where $\mu^{*}$ is a target probability supported on $\mathbb{R}^{d}$ satisfying $\mathbb{E}[\|x_{1}\|_{2}^{2}]=\int_{\mathbb{R}^{d}}\|x\|_{2}^{2}\mu^{*}({\rm d}x)<\infty$. The linear stochastic interpolant with scalar schedule is the stochastic process $I_{t}=\alpha_{t}z+\beta_{t}x_{1}$, where $z\sim\mathsf{N}(0,\mathrm{I})$ is multivariate normal distributed with $z\perp x_{1}$. Here $\alpha_{t},\beta_{t}\in C^{1}([0,1])$ are scalar functions of $t$ satisfying the boundary conditions $\alpha_{0}=\beta_{1}=1$ and $\alpha_{1}=\beta_{0}=0$, so that $I_{0}=z$ and $I_{1}=x_{1}$.
For different values of $t$, the interpolant $I_{t}$ can be seen as modeling a corruption of the target at a specific scale. The theory of stochastic interpolants [2, 1] shows that one can generate samples from $\mu^{*}$ by solving the following ODE:
(1.1) $\mathrm{d}X_{t}=b_{t}(X_{t})\mathrm{d}t,\quad X_{0}\sim\mathsf{N}(0,\mathrm{I})\,$,
where $b_{t}(x)=\mathbb{E}[\dot{I}_{t}|I_{t}=x]$ and $\dot{I}_{t}$ denotes the time derivative of $I_{t}$. The solution satisfies $\mathrm{Law}(X_{t})=\mathrm{Law}(I_{t})$, and in particular, $X_{1}\sim\mu^{*}$. This can also be seen as a consequence of the mimicking theorem [17], also referred to as Markovian projection.
Because the drift $b_{t}$ is a conditional expectation, we can define it as the minimizer of the square loss function
$L(\hat{b})=\int_{0}^{1}\mathbb{E}[\|\hat{b}_{t}(I_{t})-\dot{I}_{t}\|_{2}^{2}]\,{\rm d}t\,.$
By parametrizing $\hat{b}$ in an expressive class, using e.g. a deep neural network, and optimizing the loss function (with expectation over empirical samples), we obtain an approximation $\hat{b}\approx b$. This allows us to solve (1.1) with $\hat{b}_{t}$ to generate samples. More technical details and variants for SDEs and the a posteriori tuning of diffusion coefficients are presented in Section 2.1.
1.3. This work
Since $\hat{b}_{t}$ is learned from samples and generation requires numerical integration of a differential equation, a natural question arises: does there exist a particular choice of $\alpha_{t},\beta_{t}$ that can enhance both statistical and numerical efficiency? This paper establishes design principles for addressing this question. Specifically, our contributions are as follows:
  • In Section 2, we prove that under the Kullback-Leibler divergence criterion in path space, different choices of scalar schedules are statistically equivalent when diffusion coefficients are optimized a posteriori. This equivalence paradoxically renders statistical considerations insufficient for schedule selection.
  • In Sections 3.1 and 3.2, we introduce a principled approach to numerical optimization by minimizing the averaged squared Lipschitzness of the drift $b_{t}$ at inference time. This is enabled by a transfer formula that converts estimated drifts between interpolation schedules without retraining.
  • In Sections 3.3-3.5, we conduct analytical studies of the optimized schedules for Gaussian and Gaussian mixture distributions. For Gaussians, optimized schedules achieve exponential improvements in the Lipschitz constant; for Gaussian mixtures, they reduce mode collapse. We extend these results to log-concave and general distributions.
  • In Section 4, we demonstrate the practical benefits of optimized schedules through numerical experiments on high-dimensional Gaussian distributions and mixtures. Using insights from Gaussian analysis, we design schedules for invariant distributions of the stochastic Allen-Cahn equation and the Navier-Stokes equations, achieving improved energy spectrum estimation that remains robust across resolutions with fixed integration steps, unlike standard linear schedules.
1.4. Related work
Since the introduction of flow and diffusion models, numerous studies have examined the design principles and parameter space of these models (see a review in [48]). These investigations encompass the choice of noise, noising processes, time reversal processes, training losses, and diffusion coefficients in the generative process. This paper focuses on designing interpolation schedules within the framework of unit-time stochastic interpolants [2, 1], which relates to the noising and denoising schedules in diffusion models.
Schedule design is of interest from both statistical and numerical perspectives. From a statistical standpoint, it was demonstrated in [25] that different noise schedules in diffusion models yield the same variational lower bound. Our results suggest that this “statistical equivalence” generalizes to a broader context using the unit-time stochastic interpolants framework, with the Kullback-Leibler divergence in path space serving as the statistical estimation criterion (see discussion in Remark 2.6).
From a numerical perspective, existing works have derived insights primarily through empirical studies on machine learning datasets to tune the noise schedules for efficient sampling performance [35, 22, 30, 41, 23]; see also [36, 47, 32, 5] for learning improved schedules with additional training and an analysis [45] considering score errors. In this work, we propose a principled way for numerical design by optimizing the Lipschitzness of the drift field at inference time. Related mathematical work focused on the Lipschitz regularity of flows and flow maps includes [9, 44]. See also [3] for a mathematical investigation of the numerical impact of schedules on identifying modes in high dimensions.
It has been advocated to learn the optimal transport path [28], which is straight and therefore offers better numerical performance; see also generative models built using entropy-regularized optimal transport, namely Schrödinger bridges [11, 37]. Nevertheless, the optimal transport path may lead to irregular drift fields [44] that are not ideal for numerical integration (see also an example in Remark 3.4), an issue our proposed criterion of optimizing the Lipschitzness aims to address.
Moreover, there has been a line of work on improving numerical performance with high order, exponential, or parallel integrators, e.g., [13, 29, 50, 26, 6, 46, 10, 43], and multiscale and cascading approach [49, 12, 21, 33, 20, 16, 31], which can be combined with the design of schedules to accelerate sampling. We also note another line of work focused on consistency models and learning flow maps (e.g., [39, 24, 34, 14, 4]), to achieve few steps sampling and thus improve numerical efficiency.
2. Statistical Equivalence under Kullback-Leibler in Path Space
In this section, we discuss the statistical properties of different interpolation schedules, using the Kullback-Leibler (KL) divergence in path space as the criterion. The focus is on formal derivations and calculations, and the goal is to reveal the underlying structures rather than provide a fully rigorous treatment, which would require delicate discussions on the regularity of the SDEs.
2.1. Stochastic interpolants
Here we briefly recall the main results of the stochastic interpolant framework [2, 1]. For completeness, we also include a simple sketch of derivations in Appendix A.
As in Section 1.2, we denote the target distribution by $\mu^{*}$, and assume that it is supported on $\mathbb{R}^{d}$ and satisfies $\mathbb{E}[\|x_{1}\|_{2}^{2}]<\infty$. For simplicity we also assume that $\mu^{*}$ is absolutely continuous with respect to the Lebesgue measure and has a smooth density.
Definition 2.1.
The linear stochastic interpolant between $x_{1}\sim\mu^{*}$ and the Gaussian noise $z\sim\mathsf{N}(0,\mathrm{I})$ with $z\perp x_{1}$ is the process
(2.1) $I_{t}=\alpha_{t}z+\beta_{t}x_{1},\quad 0\leq t\leq 1\,.$
where $\alpha_{t},\beta_{t}\in C^{1}([0,1])$ are scalar interpolation schedules satisfying the boundary conditions $\alpha_{0}=\beta_{1}=1$ and $\alpha_{1}=\beta_{0}=0$ as well as $\dot{\beta}_{t}>0,\dot{\alpha}_{t}<0$ for $t\in(0,1)$.
The law of the stochastic interpolant coincide with the law of the solution of an ODE with a drift given by a conditional expectation:
Proposition 2.2.
Let $b_{t}(x)=\mathbb{E}[\dot{I}_{t}|I_{t}=x]$. Then the solutions to the ODE
${\rm d}X_{t}=b_{t}(X_{t}){\rm d}t,\quad X_{0}\sim\mathsf{N}(0,\mathrm{I})\,$,
satisfy $\mathrm{Law}(X_{t})=\mathrm{Law}(I_{t})$ for all $t\in[0,1]$, and in particular, $X_{1}\sim\mu^{*}$.
Using the Fokker-Planck equation and the fact that $\nabla\cdot(\rho\nabla\log\rho)=\Delta\rho$, we can also construct a family of SDEs that share the same law at each time as the interpolation process $I_{t}$:
Proposition 2.3.
Let $b_{t}(x)=\mathbb{E}[\dot{I}_{t}|I_{t}=x]$ and assume the density of $I_{t}$, denoted by $\rho_{t}$, exists and is $C^{1}$ in space. Then for any $\epsilon_{t}\geq 0$, the solutions to the SDE
${\rm d}X_{t}=\left(b_{t}(X_{t})+\epsilon_{t}\nabla\log\rho_{t}(X_{t})\right){\rm d}t+\sqrt{2\epsilon_{t}}{\rm d}W_{t},\quad X_{0}\sim\mathsf{N}(0,\mathrm{I})\,.$
satisfy $\mathrm{Law}(X_{t})=\mathrm{Law}(I_{t})$ for all $t\in[0,1]$, and in particular, $X_{1}\sim\mu^{*}$.
By Stein’s identity, the score $\nabla\log\rho_{t}(x)$ can be expressed as:
(2.2) $\nabla\log\rho_{t}(x)=-\frac{1}{\alpha_{t}}\mathbb{E}[z|I_{t}=x]\,.$
By using
(2.3) $x$ $=\mathbb{E}[I_{t}|I_{t}=x]=\alpha_{t}\mathbb{E}[x_{0}|I_{t}=x]+\beta_{t}\mathbb{E}[x_{1}|I_{t}=x]$
$b_{t}(x)$ $=\mathbb{E}[\dot{I}_{t}|I_{t}=x]=\dot{\alpha}_{t}\mathbb{E}[x_{0}|I_{t}=x]+\dot{\beta}_{t}\mathbb{E}[x_{1}|I_{t}=x]\,$,
after some simple algebra we can relate $b_{t}(x)$ and $\nabla\log\rho_{t}(x)$ through an affine transformation
(2.4) $b_{t}(x)$ $=\frac{\dot{\beta}_{t}}{\beta_{t}}x+\alpha_{t}^{2}(\frac{\dot{\beta}_{t}}{\beta_{t}}-\frac{\dot{\alpha}_{t}}{\alpha_{t}})\nabla\log\rho_{t}(x)\,.$
This means that, if we know $b_{t}$ or an approximation of it, we can use the above relation to obtain the score or an approximation of it directly.
2.2. Learning the drift from data
We can use empirical risk minimization to learn the conditional expectation $b$ through optimizing the square loss function
$L(\hat{b})=\int_{0}^{1}\mathbb{E}[\|\hat{b}_{t}(I_{t})-\dot{I}_{t}\|^{2}_{2}]\,{\rm d}t\,.$
In practice, the expectation is over empirical samples. Optimizing it leads to an estimate of $\hat{b}$.
It is also common to optimize for the denoiser $\mathbb{E}[x_{1}|I_{t}=x]$, or the score $\nabla\log\rho_{t}(x)=-\mathbb{E}[\frac{z}{\alpha_{t}}|I_{t}=x]$ directly. The corresponding loss functions can be similarly constructed since these terms are all expressed as conditional expectations. We note that the three objects can be recovered from each other by affine transformations, using (2.3) and (2.4). Thus, without loss of generality and for a unified analysis, let us assume that at the end we have an estimator of the score in terms of $\hat{s}_{t}(x)\approx\nabla\log\rho_{t}(x)$. This means that the estimated SDE has the form
${\rm d}\hat{X}_{t}=\left(\frac{\dot{\beta}_{t}}{\beta_{t}}\hat{X}_{t}+(\alpha_{t}^{2}(\frac{\dot{\beta}_{t}}{\beta_{t}}-\frac{\dot{\alpha}_{t}}{\alpha_{t}})+\epsilon_{t})\hat{s}_{t}(x)\right){\rm d}t+\sqrt{2\epsilon_{t}}{\rm d}W_{t},\quad\hat{X}_{0}\sim\mathsf{N}(0,\mathrm{I})\,.$
2.3. Optimizing the KL in path space
Given the flexibility of choosing $\epsilon_{t}$, it is natural to ask which $\epsilon_{t}$ is optimal. Let us consider the criterion of the KL divergence between path measures $\mathbb{P}_{X}$ and $\mathbb{P}_{\hat{X}}$ of $X=(X_{t})_{0\leq t\leq 1}$ and $\hat{X}=(\hat{X}_{t})_{0\leq t\leq 1}$, respectively. Aaccording to Girsanov’s theorem, this KL divergence has the form
(2.5) $\mathrm{KL}[\mathbb{P}_{X}\|\mathbb{P}_{\hat{X}}]=\frac{1}{2\epsilon_{t}}\int_{0}^{1}\left(\alpha_{t}^{2}(\frac{\dot{\beta}_{t}}{\beta_{t}}-\frac{\dot{\alpha}_{t}}{\alpha_{t}})+\epsilon_{t}\right)^{2}\|\nabla\log\rho_{t}(x)-\hat{s}_{t}(x)\|^{2}_{2}\rho_{t}(x){\rm d}t\,.$
Now, recall the fact that, for any $a$, the minimizer of $\frac{(\epsilon+a)^{2}}{2\epsilon}=\frac{\epsilon}{2}+a+\frac{a^{2}}{2\epsilon}$ is $\epsilon=|a|$, and the minimum is $\max\{0,2a\}$. Thus, the KL achieves minimum when $\epsilon_{t}=\alpha_{t}^{2}(\frac{\dot{\beta}_{t}}{\beta_{t}}-\frac{\dot{\alpha}_{t}}{\alpha_{t}})$. Viewing this optimized KL as a function of the interpolation schedules $\alpha,\beta$ and denoting it as $\mathrm{KL}^{\star}(\alpha,\beta)$, it reads
(2.6) $\mathrm{KL}^{\star}(\alpha,\beta)=2\int_{\mathbb{R}^{d}}\int_{0}^{1}\alpha_{t}^{2}(\frac{\dot{\beta}_{t}}{\beta_{t}}-\frac{\dot{\alpha}_{t}}{\alpha_{t}})\|\nabla\log\rho_{t}(x)-\hat{s}_{t}(x)\|^{2}_{2}\rho_{t}(x){\rm d}t{\rm d}x\,.$
Remark 2.4.
For certain choices of $\alpha_{t},\beta_{t}$, the resulting $\epsilon_{t}$ may blow up. However, the SDE is still well defined; see examples in Appendix B.  $\lozenge$
2.4. Equivalence over scalar schedules
Our next result shows that, remarkably, $\mathrm{KL}^{\star}(\alpha,\beta)$ remains constant regardless of the interpolation schedules $\alpha_{t},\beta_{t}$ we choose.
Proposition 2.5.
Let $q_{\eta}(x)$ be the probability density function of $x_{1}+\eta z$ with $\eta\geq 0$ and denote by $\hat{S}_{\eta}(x)$ the estimator of its score $\nabla\log q_{\eta}(x)$. Then
(2.7) $\mathrm{KL}^{\star}(\alpha,\beta)=2\int_{0}^{\infty}\eta\cdot\mathbb{E}[\|\nabla\log q_{r}(x_{1}+\eta z)-\hat{S}_{r}(x_{1}+\eta z)\|^{2}_{2}]{\rm d}\eta\,.$
Proof.
We know that $\rho_{t}(x)$ is the density of $\alpha_{t}z+\beta_{t}x_{1}=\beta_{t}(x_{1}+\frac{\alpha_{t}}{\beta_{t}}z)$. Thus $\nabla\log\rho_{t}(x)=\frac{1}{\beta_{t}}\nabla\log q_{\frac{\alpha_{t}}{\beta_{t}}}(\frac{x}{\beta_{t}})$, and $\hat{s}_{t}(x)=\frac{1}{\beta_{t}}\hat{S}_{\frac{\alpha_{t}}{\beta_{t}}}(\frac{x}{\beta_{t}})$ where $\hat{S}_{\eta}(x)=\nabla\log q_{\eta}(x)$. Using these relations, we have
(2.8) $\mathrm{KL}^{\star}(\alpha,\beta)$ $=2\int_{\mathbb{R}^{d}}\int_{0}^{1}\frac{\alpha_{t}^{2}}{\beta_{t}^{2}}(\frac{\dot{\beta}_{t}}{\beta_{t}}-\frac{\dot{\alpha}_{t}}{\alpha_{t}})\|\nabla\log q_{\frac{\alpha_{t}}{\beta_{t}}}(\frac{x}{\beta_{t}})-\hat{S}_{\frac{\alpha_{t}}{\beta_{t}}}(\frac{x}{\beta_{t}})\|^{2}_{2}\rho_{t}(x){\rm d}t{\rm d}x$
$=2\int_{0}^{1}\frac{\alpha_{t}^{2}}{\beta_{t}^{2}}(\frac{\dot{\beta}_{t}}{\beta_{t}}-\frac{\dot{\alpha}_{t}}{\alpha_{t}})\mathbb{E}[\|\nabla\log q_{\frac{\alpha_{t}}{\beta_{t}}}(x_{1}+\frac{\alpha_{t}}{\beta_{t}}z)-\hat{S}_{\frac{\alpha_{t}}{\beta_{t}}}(x_{1}+\frac{\alpha_{t}}{\beta_{t}}z)\|^{2}_{2}]{\rm d}t\,.$
Noting that $\frac{\alpha_{t}^{2}}{\beta_{t}^{2}}(\frac{\dot{\beta}_{t}}{\beta_{t}}-\frac{\dot{\alpha}_{t}}{\alpha_{t}})=-\frac{\alpha_{t}}{\beta_{t}}\frac{\rm d}{{\rm d}t}(\frac{\alpha_{t}}{\beta_{t}})$ and using $\alpha_{t}/\beta_{t}$ instead of $t$ as integration variable, we arrive at (2.7). ∎
Remark 2.6.
In [25], it was pointed out that in diffusion models, different noise schedules lead to the same variational lower bound. In the continuous setting, this corresponds to the KL divergence in path space. Our results generalize their discussion to stochastic interpolants and incorporate the step of a posteriori tuning of diffusion coefficients.  $\lozenge$
Proposition 2.5 shows that the optimal KL accuracy in path space depends solely on the estimation of $\nabla\log q_{r}(x_{1}+rz)$: that is, from the perspective of KL divergence in path space, all linear scalar interpolants with independently coupled endpoints and one endpoint Gaussian are statistically indistinguishable. This indicates that other metrics need to be explored if we want to select models for improved statistical efficiency. On the other hand, using matrix-valued instead of scalar schedules may potentially lead to different statistical efficiency, a direction of interest in future work.
3. Numerical Design by Optimizing Averaged Squared Lipschitzness
The discussion in the previous section does not consider numerical efficiency: while different scalar schedules are statistically equivalent, they lead to ODEs or SDEs with dramatically different regularity properties of the drift term. In this section, we explore how to choose interpolation schedules that enhance numerical efficiency. We focus on ODEs rather than SDEs for simplicity, noting that ODEs typically achieve better empirical performance due to their greater ease of integration [23, 13].
3.1. From one schedule to another
First, we point out a fact that given the estimated drift for one particular scalar interpolation schedule, one can directly obtain an estimated drift at another arbitrary scalar interpolation schedule. Without loss of generality, we consider one reference scalar schedule $\alpha^{\dagger}_{t}=1-t,\beta^{\dagger}_{t}=t$.
Proposition 3.1.
Consider the two stochastic interpolants $I_{t}^{\dagger}=\alpha^{\dagger}_{t}z+\beta^{\dagger}_{t}x_{1}$ and $I_{t}=\alpha_{t}z+\beta_{t}x_{1}$ and their associated drifts $b^{\dagger}(x)=\mathbb{E}[\dot{I}_{t}^{\dagger}|I_{t}=x]$ and $b_{t}(x)=\mathbb{E}[\dot{I}_{t}|I_{t}=x]$. Then with $t^{\dagger}=1/(1+\alpha_{t}/\beta_{t})$, it holds that
(3.1) $b_{t}(x)=\frac{\dot{\alpha}_{t}}{\alpha_{t}}x+\Big{(}\dot{\beta}_{t}-\frac{\dot{\alpha}_{t}\beta_{t}}{\alpha_{t}}\Big{)}\left((1-t^{\dagger})b^{\dagger}_{t^{\dagger}}\Big{(}\frac{t^{\dagger}}{\beta_{t}}x\Big{)}+\frac{t^{\dagger}}{\beta_{t}}x\right)\,.$
Proof.
By direct algebraic calculations, we get
(3.2) $b^{\dagger}_{t}(x)$ $=\mathbb{E}[x_{1}-z|I_{t}=x]=\mathbb{E}[x_{1}-\frac{I_{t}-tx_{1}}{1-t}|I_{t}=x]$
$=-\frac{x}{1-t}+\frac{1}{1-t}\mathbb{E}[x_{1}|x_{1}+\frac{1-t}{t}z=\frac{x}{t}]\,$,
and similarly
(3.3) $b_{t}(x)$ $=\mathbb{E}[\dot{\alpha}_{t}z+\dot{\beta}_{t}x_{1}|I_{t}=x]=\mathbb{E}[\dot{\alpha}_{t}\frac{I_{t}-\beta_{t}x_{1}}{\alpha_{t}}+\dot{\beta}_{t}x_{1}|I_{t}=x]$
$=\frac{\dot{\alpha}_{t}}{\alpha_{t}}x+(\dot{\beta}_{t}-\frac{\dot{\alpha}_{t}\beta_{t}}{\alpha_{t}})\mathbb{E}[x_{1}|x_{1}+\frac{\alpha_{t}}{\beta_{t}}z=\frac{x}{\beta_{t}}]\,.$
Let $t^{\dagger}$ satisfy $\alpha_{t}/\beta_{t}=(1-t^{\dagger})/t^{\dagger}$. This means that $t^{\dagger}=1/(1+\alpha_{t}/\beta_{t})$. Therefore, combining (3.2) and (3.3), we arrive at (3.1). ∎
The proposition implies that we can easily change the interpolation schedule from one to another if we know the true drift functions. This also applies to the estimators of the drift functions, so we can tune the schedule at inference time rather than during training. Similar statements have appeared in the literature [25, 23]. We will use this fact in numerical experiments in Section 4. The natural question now is which schedule to choose in practice.
3.2. Optimizing averaged squared Lipschitzness
As natural and principled approach to choose the schedule, we propose to minimize the following averaged squared Lipschitzness criterion.
Definition 3.2.
The averaged squared Lipschitzness (avg-Lip2) is defined as
(3.4) $A_{2}=\int_{0}^{1}\mathbb{E}[\|\nabla b_{t}(I_{t})\|^{2}_{2}]\,\mathrm{d}t\,$,
where $\|\cdot\|_{2}$ is the $2$-norm.
In general, we could optimize $A_{2}$ over all possible nonlinear interpolants $I_{t}$. Here, for simplicity, we restrict ourselves to linear interpolants with scalar schedules $\alpha,\beta$ [^1] See discussions on matrix-valued schedules in Remark 3.10.. We provide several examples in the next two sections and show the significance of this criterion in numerical performance and compares it with optimal transport.
3.3. 1D example: Gaussian
We begin with analytic studies on 1D Gaussians.
Example 3.3 (1D Gaussian).
Consider $I_{t}=\alpha_{t}z+\beta_{t}x_{1}$ with $x_{1}\sim\mathsf{N}(0,M)\perp z\sim\mathsf{N}(0,1)$. Here $M>0$ is a positive scalar. Then
$b_{t}(x)=\mathbb{E}[\dot{I}_{t}|I_{t}=x]=\mathrm{Cov}(\dot{I}_{t},I_{t})\mathrm{Cov}(I_{t})^{-1}x=(\alpha_{t}\dot{\alpha}_{t}+\beta_{t}\dot{\beta}_{t}M)(\alpha_{t}^{2}+\beta_{t}^{2}M)^{-1}x\,.$
If we take $\alpha_{t}=1-t,\beta_{t}=t$, we get
$b_{t}(x)=\frac{t-1+tM}{(1-t)^{2}+t^{2}M}x\,.$
Suppose $M$ is a large number [^2] Although we can always use variance preserving design to fix this setting, it may still occur for a particular Fourier frequency component in high high-dimensional setting. Similar discussions apply when $M$ is a small number.. We have
$A_{2}=\int_{0}^{1}\frac{(t-1+tM)^{2}}{((1-t)^{2}+t^{2}M)^{2}}{\rm d}t\geq\int_{\frac{1}{M^{1/3}}}^{\frac{1}{M^{1/2}}}\frac{(t-1+tM)^{2}}{((1-t)^{2}+t^{2}M)^{2}}{\rm d}t\geq\Omega(\sqrt{M})\,.$
Moreover, the Lipschitzness $\|\nabla b_{t}(1/M)\|_{2}\geq\Omega(M)$ which grows linearly with $M$.
However, we can optimize
(3.5) $A_{2}=\int_{0}^{1}\mathbb{E}[\|\nabla b_{t}(I_{t})\|_{2}^{2}]\,{\rm d}t$ $=\int_{0}^{1}\mathbb{E}[\|\mathrm{Cov}(\dot{I}_{t},I_{t})\mathrm{Cov}(I_{t})^{-1}\|_{2}^{2}]\,{\rm d}t$
$=\frac{1}{4}\int_{0}^{1}\left\|\frac{{\rm d}}{{\rm d}t}\log\mathrm{Cov}(I_{t})\right\|_{2}^{2}{\rm d}t\,.$
By Cauchy–Schwarz inequality, the minimizer satisfies $\frac{{\rm d}}{{\rm d}t}\log\mathrm{Cov}(I_{t})=\mathrm{const}$. To achieve the minimum, we get $\log\mathrm{Cov}(I_{t})=(1-t)\log\mathrm{Cov}(I_{0})+t\log\mathrm{Cov}(I_{1})$. Solving this equation yields $\alpha_{t}^{2}+\beta_{t}^{2}M=M^{t}$. Taking the choice $\alpha_{t}^{2}=1-\beta^{2}_{t}$, we obtain the interpolation schedule
(3.6) $\alpha_{t}=\sqrt{\frac{M-M^{t}}{M-1}},\beta_{t}=\sqrt{\frac{M^{t}-1}{M-1}}\,.$
For such choice, $b_{t}(x)=\frac{1}{2}(\log M)x$. The corresponding $A_{2}=O(\log^{2}M)$ and $\|\nabla b_{t}(x)\|_{2}\leq\frac{1}{2}|\log M|$ for all $t\in[0,1],x\in\mathbb{R}$. This shows that there is an exponential improvement in the averaged squared Lipschitzness and the actual Lipschitz constant of the drift, compared to $\alpha_{t}=1-t,\beta_{t}=t$.
Remark 3.4.
We compare the above to optimal transport, which minimizes the squared path length $P=\int_{0}^{1}\mathbb{E}[\|b_{t}(I_{t})\|_{2}^{2}]\,{\rm d}t$. Using the optimal transport theory [^3] See details in Appendix C.1., we get that
$b_{t}(x)=\frac{\sqrt{M}-1}{1-t+t\sqrt{M}}x\,.$
This can have a large Lipschitz constant near $t=0$ when $M$ is large.  $\lozenge$
3.4. 1D example: Gaussian mixture
We then move to Gaussian mixture.
Example 3.5 (1D Gaussian mixtures).
Consider the 1D bimodal Gaussian mixture
$\mu^{*}(x)=p\mathsf{N}(x;M,1)+(1-p)\mathsf{N}(x;-M,1)\,.$
To enable an explicit analytic study [^4] See calculation details in Remark C.3 in Appendix C., we take $\alpha_{t}=\sqrt{1-\beta_{t}^{2}}$, which leads to
(3.7) $b_{t}(x)=\dot{\beta}_{t}M\mathrm{tanh}(h+\beta_{t}Mx)\,$,
where $h$ satisfies $\frac{p}{1-p}=\exp(2h)$, or equivalently $p=\frac{\exp(h)}{\exp(h)+\exp(-h)}$.
Suppose $h>0$. If $\beta_{t}=t$ and $M$ is large, we observe that at the initial time, $b_{0}(x)=M\tanh(h)$, which is large. In the one-dimensional case, this means all points move toward the right when using a forward Euler discretization with step size $O(1)$. Even for negative $x$, such a drift will likely push these points into positive territory. On the other hand, we know that for $x>0$, we have $b_{t}(x)>0$. This means that once a point reaches the positive side, it will remain positive. Therefore, such a discretization scheme will miss the mode on the left side. The above argument demonstrates that we must use an initial step size of $O(1/M)$ to ensure that the discretization does not miss modes.
Below, we study the optimization of avg-Lip2, which leads to a schedule $\beta$ that grows slowly at initial time that does not suffer from the mode missing issue, namely, we can safely use a discretization scheme with uniform stepsize.
Proposition 3.6 (Optimizing avg-Lip2 for 1D Gaussian mixture).
For the 1D bimodal Gaussian mixture example, if we optimize $A_{2}$ over all possible linear interpolants $I_{t}$ with scalar schedules satisfying $\alpha_{t}^{2}+\beta_{t}^{2}=1$, then the optimal $\beta_{t}$ satisfies $(0\leq t\ \leq 1)$
(3.8) $t=\frac{\int_{0}^{\beta_{t}}u(G(u))^{1/2}{\rm d}u}{\int_{0}^{1}u(G(u))^{1/2}{\rm d}u}\,$,
where $G(u)=\mathbb{E}[\operatorname{sech}^{4}(h+uM(\sqrt{1-u^{2}}z+ux_{1}))]$. Equivalently, we have the following Euler-Lagrange equation for the optimal $\beta_{t}$:
$\dot{\beta}_{t}^{2}\beta_{t}-\ddot{\beta}_{t}\beta_{t}^{2}+2\dot{\beta}_{t}^{2}\beta_{t}^{3}M^{2}(1+\frac{3}{4}\operatorname{Corr}(I_{t}\operatorname{tanh}(h+\beta_{t}MI_{t}),\operatorname{sech}^{4}(h+\beta_{t}MI_{t})))=0\,$,
where $I_{t}=\sqrt{1-\beta^{2}_{t}}z+\beta_{t}x_{1}$. If we omit the Corr term, we get $\dot{\beta}_{t}^{2}\beta_{t}-\ddot{\beta}_{t}\beta_{t}^{2}+2\dot{\beta}_{t}^{2}\beta_{t}^{2}M^{2}=0$ which has the solution
(3.9) $\beta_{t}=\frac{1}{M}\sqrt{-\log(1+(e^{-M^{2}}-1)t)}\,.$
The proof of this proposition is in Appendix C.3.
Remark 3.7.
The time-dilated schedule studied in [3] also resolve the mode missing issue:
(3.10) $\beta_{t}=\begin{cases}\displaystyle\frac{2\kappa\,t}{M},&t\in\bigl{[}0,\tfrac{1}{2}\bigr{]}\,,\\[4.30554pt] \displaystyle\frac{\kappa}{M}+\Bigl{(}1-\frac{\kappa}{M}\Bigr{)}\,(2t-1),&t\in\bigl{[}\tfrac{1}{2},1\bigr{]}\,.\end{cases}$
where $\kappa$ is a constant.  $\lozenge$
We plot different schedules in Figure 1 and we solve for the true solutions numerically using (3.8). The dilated (3.10), optimal min-avg-Lip2 (3.8), and approximate min-avg-Lip2 solution (3.9) all exhibit slower growth near $t=0$ compared to the standard linear schedule. Their key difference lies in their behavior near $t=1$. The optimal and approximate min-avg-Lip2 solutions exhibit more rapid growth near $t=1$, which may cause numerical issues. However, their initial slowness allows the method to sample both modes without using a small stepsize, as we demonstrate in Section 4.2.
Figure 1. Comparison of different interpolation schedules $\beta_{t}$. Left: $M=5$. Right: $M=20$. We set $p=0.3$. For the dilated schedule, we take $\kappa=1$.
Remark 3.8.
One may optimize instead $\int_{0}^{1}\mathbb{E}[\|\nabla b_{t}(I_{t})\|^{2k}_{2}]\,\mathrm{d}t$, then the optimal $\beta_{t}$ will satisfy
$t=\frac{\int_{0}^{\beta_{t}}u(G(u))^{1/{2k}}{\rm d}u}{\int_{0}^{1}u(G(u))^{1/{2k}}{\rm d}u}\,$,
where now $G(u)=\mathbb{E}[\operatorname{sech}^{4k}(h+uM(\sqrt{1-u^{2}}z+ux_{1}))]$ and a similar ODE for $\beta_{t}$ holds. For details, see Appendix C.3. Detailed investigation of choice of $k$ is out of the scope of this paper, which may improve the behavior near $t=1$.  $\lozenge$
3.5. High dimensional examples
We then move beyond 1D examples.
Proposition 3.9 (Optimizing avg-Lip2 for high dimensional Gaussians).
Consider $x_{1}\sim\mathsf{N}(0,M)\perp z\sim\mathsf{N}(0,\mathrm{I})$ in $d$ dimensions with $M$ now a positive-definite symmetric matrix. Denote the eigendecomposition $M=U\Lambda U^{T}$ where $U$ is an orthogonal matrix and $\Lambda=\mathrm{diag}(\lambda^{(1)},...,\lambda^{(d)})$ with $1\geq\lambda^{(1)}\geq\lambda^{(2)}\geq...\geq\lambda^{(d)}>0$.
If we optimize $A_{2}$ over all possible linear interpolants $I_{t}$ with scalar schedules, then, the optimal solution is $I_{t}=\alpha_{t}z+\beta_{t}x_{1}$ with
(3.11) $\alpha_{t}=\sqrt{\frac{\lambda^{\star}-(\lambda^{\star})^{t}}{\lambda^{\star}-1}},\beta_{t}=\sqrt{\frac{(\lambda^{\star})^{t}-1}{\lambda^{\star}-1}}\,.$
where $\lambda^{\star}=\lambda^{(d)}$. For the optimal solution, the corresponding 2-norm $\|\nabla b_{t}(x)\|_{2}=\frac{1}{2}|\log\lambda^{\star}|$.
Proof of Proposition 3.9.
First, because the interpolant is linear and $z,x_{1}$ are jointly Gaussian, we have that $I_{t},\dot{I}_{t}$ are jointly Gaussian. Thus,
$b_{t}(x)=\mathbb{E}[\dot{I}_{t}|I_{t}=x]=\mathrm{Cov}(\dot{I}_{t},I_{t})\mathrm{Cov}(I_{t})^{-1}x=(\alpha_{t}\dot{\alpha}_{t}+\beta_{t}\dot{\beta}_{t}M)(\alpha_{t}^{2}+\beta_{t}^{2}M)^{-1}x\,.$
We can calculate the $2$-norm using the eigenvalues:
$\|\nabla b_{t}(x)\|_{2}=\max_{1\leq j\leq d}\left|\frac{\alpha_{t}\dot{\alpha}_{t}+\beta_{t}\dot{\beta}_{t}\lambda^{(j)}}{\alpha_{t}^{2}+\beta_{t}^{2}\lambda^{(j)}}\right|=\max\{\left|\frac{\alpha_{t}\dot{\alpha}_{t}+\beta_{t}\dot{\beta}_{t}\lambda^{(1)}}{\alpha_{t}^{2}+\beta_{t}^{2}\lambda^{(1)}}\right|,\left|\frac{\alpha_{t}\dot{\alpha}_{t}+\beta_{t}\dot{\beta}_{t}\lambda^{(d)}}{\alpha_{t}^{2}+\beta_{t}^{2}\lambda^{(d)}}\right|\}\,$,
where, in the last equality, we used the fact that the function $\lambda\to\frac{\alpha_{t}\dot{\alpha}_{t}+\beta_{t}\dot{\beta}_{t}\lambda}{\alpha_{t}^{2}+\beta_{t}^{2}\lambda}$ is non-decreasing. This implies that for $\lambda=\lambda^{(1)}$ or $\lambda^{(d)}$,
$A_{2}=\int_{0}^{1}\mathbb{E}[\|\nabla b_{t}(I_{t})\|^{2}_{2}]{\rm d}t\geq\int_{0}^{1}\left|\frac{\alpha_{t}\dot{\alpha}_{t}+\beta_{t}\dot{\beta}_{t}\lambda}{\alpha_{t}^{2}+\beta_{t}^{2}\lambda}\right|^{2}{\rm d}t=\frac{1}{4}\int_{0}^{1}\left|\frac{{\rm d}}{{\rm d}t}\log(\alpha_{t}^{2}+\beta_{t}^{2}\lambda)\right|^{2}{\rm d}t\,.$
By Cauchy–Schwarz inequality, $A_{2}\geq\frac{1}{4}\log^{2}\lambda$ for $\lambda=\lambda^{(1)}$ or $\lambda^{(d)}$. Using the assumption and definition $\lambda^{\star}$, we have $A_{2}\geq\frac{1}{4}\log^{2}\lambda^{\star}$. Similar to the discussion in Section 3.3, the minimum can be achieved by taking $\frac{{\rm d}}{{\rm d}t}\log(\alpha_{t}^{2}+\beta_{t}^{2}\lambda^{\star})=\log\lambda^{\star}$; the assumption $1\geq\lambda^{(1)}$ is used to verify the minimum. Taking $\alpha_{t}=\sqrt{1-\beta^{2}_{t}}$ then leads to the solution in (3.11). ∎
Proposition 3.9 shows that by adapting the interpolation schedules, the Lipschitz constant of the drift field depends on the magnitude of eigenvalues logarithmically, compared to algebraically when using the simple schedule $\alpha_{t}=1-t,\beta_{t}=t$. This is similar to the discussion for the 1D case in Section 3.3.
Remark 3.10 (Discussions on matrix-valued schedules).
If we allow matrix-valued schedules, it is possible to further improve numerical efficiency by adapting the schedule to each eigenvalue individually. In detail, consider the following choice:
$\alpha_{t}=U\mathrm{diag}(\alpha^{(1)}_{t},\ldots,\alpha^{(d)}_{t})U^{T},\quad\beta_{t}=U\mathrm{diag}(\beta^{(1)}_{t},\ldots,\beta^{(d)}_{t})U^{T}\,$,
where
$\alpha_{t}^{(k)}=\sqrt{\frac{\lambda^{(k)}-(\lambda^{(k)})^{t}}{\lambda^{(k)}-1}},\quad\beta_{t}^{(k)}=\sqrt{\frac{(\lambda^{(k)})^{t}-1}{\lambda^{(k)}-1}}\,.$
When $\lambda^{(k)}=1$, we interpret this formula through the limit $\lambda^{(k)}\to 1$. Direct calculation using this formula yields
$b_{t}(x)=\mathrm{Cov}(\dot{I}_{t},I_{t})\mathrm{Cov}(I_{t})^{-1}x$ $=(\dot{\alpha}_{t}\alpha_{t}^{T}+\dot{\beta}_{t}M\beta_{t}^{T})(\alpha_{t}\alpha_{t}^{T}+\beta_{t}M\beta_{t}^{T})^{-1}x$
$=\frac{1}{2}U\mathrm{diag}(\log\lambda^{(1)},\ldots,\log\lambda^{(d)})U^{T}x\,.$
Here, each eigenvector direction corresponds to its individual Lipschitz constant $|\log\lambda^{(i)}|$ for $1\leq i\leq d$, and not all scales suffer from the largest $|\log\lambda^{\star}|$. We leave the investigation of matrix-valued schedules for future study.  $\lozenge$
Example 3.11 (Extension to log-concave distributions).
We can generalize the discussion of high-dimensional Gaussians to log-concave distributions. Let $\mu^{*}\propto\exp(-V)$ with $V\in C^{2}(\mathbb{R}^{d})$ and $\lambda_{m}I\preceq\nabla^{2}V\preceq\lambda_{M}I$ where we assume $\lambda_{m}\geq 1$. Consider $x_{1}\sim\mu^{*}$ independent of $z\sim\mathsf{N}(0,\mathrm{I})$. Then for the linear interpolant with scalar schedule $I_{t}=\alpha_{t}z+\beta_{t}x_{1}$, we have
$\frac{\alpha_{t}\dot{\alpha}_{t}+\beta_{t}\dot{\beta}_{t}\lambda_{M}^{-1}}{\alpha_{t}^{2}+\beta_{t}^{2}\lambda_{M}^{-1}}\preceq\nabla b_{t}(x)\preceq\frac{\alpha_{t}\dot{\alpha}_{t}+\beta_{t}\dot{\beta}_{t}\lambda_{m}^{-1}}{\alpha_{t}^{2}+\beta_{t}^{2}\lambda_{m}^{-1}}\,.$
This can be proved using the Cramér–Rao and Brascamp–Lieb inequalities; see [15]. Therefore, similar to the Gaussian case, we can choose $\lambda^{\star}=\lambda_{M}^{-1}$. Then, with the schedule
(3.12) $\alpha_{t}=\sqrt{\frac{\lambda^{\star}-(\lambda^{\star})^{t}}{\lambda^{\star}-1}},\quad\beta_{t}=\sqrt{\frac{(\lambda^{\star})^{t}-1}{\lambda^{\star}-1}}\,$,
we have $\|\nabla b_{t}(x)\|_{2}\leq\frac{1}{2}|\log\lambda^{\star}|$. In general, we do not know an explicit solution for optimizing $A_{2}$ for log-concave distributions. However, the above schedule serves as a good choice, and the bound is tight and yields the optimal $A_{2}$ when the log-concave distribution is Gaussian.
Example 3.12 (A particular example on high dimensional Gaussian mixtures).
Consider the bimodal Gaussian mixture in $d$ dimensions
(3.13) $\mu^{*}(x)=p\mathsf{N}(x;r,\mathrm{I})+(1-p)\mathsf{N}(x;-r,\mathrm{I})\,$,
where $x\in\mathbb{R}^{d}$, and $r\in\mathbb{R}^{d}$ is a fixed vector satisfying $\|r\|_{2}=\sqrt{d}$; for instance, $r=(1,1,...,1)^{T}$. The interpolant $I_{t}=\alpha_{t}z+\beta_{t}x_{1}$ where $z\sim\mathsf{N}(0,\mathrm{I})\perp x_{1}\sim\mu^{*}$.
Using the general formula in Appendix C.2, we get $b_{t}(x)=\dot{\beta}_{t}r\mathrm{tanh}(h+\beta_{t}\langle r,x\rangle)$. Then $\nabla b_{t}(x)=\dot{\beta}_{t}\beta_{t}rr^{T}\operatorname{sech}^{2}(h+\beta_{t}\langle r,x\rangle)$, which yields
$\|\nabla b_{t}(x)\|_{2}^{2}=d\dot{\beta}_{t}^{2}\beta_{t}^{2}\operatorname{sech}^{4}(h+\beta_{t}\langle r,x\rangle)\,.$
This is effectively the same as the 1D example in Proposition 3.6. Using the result there, we get that the optimal $\beta_{t},\alpha_{t}=\sqrt{1-\beta_{t}^{2}}$ minimizing $A_{2}$ satisfies
$t=\frac{\int_{0}^{\beta_{t}}u(G(u))^{1/2}{\rm d}u}{\int_{0}^{1}u(G(u))^{1/2}{\rm d}u}\,.$
where $G(u)=\mathbb{E}[\operatorname{sech}^{4}(h+u\langle r,\sqrt{1-u^{2}}z+ux_{1}\rangle)]$. Again, an approximate solution is
(3.14) $\beta_{t}=\frac{1}{\sqrt{d}}\sqrt{-\log(1+(e^{-d}-1)t)}\,.$
Beyond the above examples, we have a general formula for optimizing $A_{2}$ over scalar interpolation schedules, for general distributions.
Example 3.13 (Optimizing avg-Lip2 for general distributions).
Consider a general distribution $\mu^{*}$ in $d$ dimensions and we assume it to be smooth for simplicity. Let $b^{\dagger}(x)$ be defined as in Proposition 3.1 and let $\alpha_{t}=\sqrt{1-\beta_{t}^{2}}$. Then using Proposition 3.1,
$b_{t}(x)=\dot{\beta}_{t}\left(\frac{-\beta_{t}}{1-\beta_{t}^{2}}x+\frac{1}{1-\beta_{t}^{2}}\left((1-t^{\dagger})b^{\dagger}_{t^{\dagger}}(\frac{t^{\dagger}}{\beta_{t}}x)+x\right)\right)\,$,
and
$\nabla b_{t}(x)=\dot{\beta}_{t}\left(\frac{-\beta_{t}}{1-\beta_{t}^{2}}\mathrm{I}+\frac{1}{1-\beta_{t}^{2}}\left((1-t^{\dagger})\frac{t^{\dagger}}{\beta_{t}}\nabla b^{\dagger}_{t^{\dagger}}(\frac{t^{\dagger}}{\beta_{t}}x)+\mathrm{I}\right)\right)=\dot{\beta}_{t}F(\beta_{t},x)\,$,
where we denote the term in the big bracket by $F(\beta_{t},x)$. Then
$A_{2}=\int_{0}^{1}\mathbb{E}[\|\nabla b_{t}(I_{t})\|_{2}^{2}]=\int_{0}^{1}\dot{\beta}_{t}^{2}\mathbb{E}[\|F(\beta_{t},I_{t})\|_{2}^{2}]{\rm d}t=\int_{0}^{1}\dot{\beta}_{t}^{2}G(\beta_{t}){\rm d}t\,$,
where we denote $G(\beta_{t})=\mathbb{E}[\|F(\beta_{t},I_{t})\|_{2}^{2}]$. Solving the Euler-Lagrange equation with the Beltrami Identity (see Appendix C.3) leads to the equation that $\beta_{t}$ satisfies:
$t=\frac{\int_{0}^{\beta_{t}}(G(u))^{1/2}{\rm d}u}{\int_{0}^{1}(G(u))^{1/2}{\rm d}u}\,.$
In general, finding the optimal $\beta_{t}$ analytically is challenging. While numerical solutions are possible once $b^{\dagger}$ is available, it is computationally costly in high dimensions as we need to evaluate $G$. Our previous examples demonstrate that certain cases allow for simpler solutions. In particular, we have an analytic formula for the Gaussian case. For Gaussian mixture distributions, we can derive approximate analytical solutions, and for log-concave cases, we can leverage insights from the Gaussian analysis to construct schedules that achieve our numerical objectives.
4. Numerical Demonstrations
In this section, we conduct numerical experiments to demonstrate the improved efficiency of the schedule designed in the previous section. We first study high-dimensional Gaussian distributions and Gaussian mixtures, followed by the high-dimensional invariant distributions of the stochastic Allen-Cahn equation and the Navier-Stokes equations, which exhibit slightly and highly non-Gaussian behaviors respectively.
We use the UNet architecture popularized by [19] to train the drift field for all experiments except Gaussian and mixtures, where we employ explicit formula. In all the examples, we integrate the ODE from $t_{\min}=10^{-3}$ to $t_{\max}=1-10^{-3}$ to avoid potential numerical issues at $t=0$ or $1$. Code is available at https://github.com/yifanc96/GenerativeDynamics-NumericalDesign.git.
For accuracy evaluation, we use the energy spectrum (or enstrophy spectrum in the case of Navier-Stokes) of the samples as the criterion. The spectrum for a sample $u$ (which is a function that is either 1D or 2D in this paper) is computed using the formula
$E(k)=\sum_{k\leq|m|_{2}\leq k+1}|\hat{u}(m)|^{2}\,$,
where $\hat{u}(m)$ are the Fourier coefficients. We average $E(k)$ over sufficiently many samples for each frequency $k$.
4.1. Gaussians
We consider the Gaussian random field $\mathsf{N}(0,\sigma^{2}(-\Delta+\tau^{2}I)^{-s})$, where $\Delta$ is the negative Laplacian with homogeneous Dirichlet boundary conditions on $D=[0,1]^{2}$. The true data distribution $x_{1}$ is sampled from this distribution with parameters $s=3$, $\tau=1$, and $\sigma^{2}=(4\pi^{2}+\tau^{2})^{s}$. The noise $z$ in the interpolant is sampled from white noise (which corresponds to $\sigma=1,s=0$ in the Gaussian random field). We discretize the 2D field on a grid with $N$ points in each dimension and construct flow-based generative models. The ODE is solved using the fourth-order Runge-Kutta (RK4) scheme.
Figure 2 shows random fields generated using a standard linear schedule $\beta_{t}=t$ compared to those using our designed schedule (3.11) optimized for avg-Lip2, both employing 20 RK4 steps with $N=128$. The designed schedule clearly produces superior samples. The right panel of Figure 2 displays both schedules: notably, the designed schedule exhibits rapid initial growth.
In Figure 3, we compare the energy spectra of the true distribution and generated samples. The designed schedule yields a more accurate spectrum, and this accuracy remains robust as resolution increases, unlike standard linear schedules where performance degrades with refinement.
Figure 2. Left: $128\times 128$ Gaussian fields generated by using linear schedules with 2020 steps of the RK4 integrator. Middle: $128\times 128$ Gaussian fields generated by using the designed schedules with 2020 steps of the RK4 integrator. Right: linear and designed schedules.
Figure 3. Energy spectra of Gaussian fields: comparison between truth, generated via designed schedules or standard linear schedules, with $20,40$ or $80$ RK4 steps. The three figures correspond to different resolutions. Left: $32\times 32$; middle: $64\times 64$; right: $128\times 128$.
4.2. Gaussian mixtures
We consider the $d$-dimensional Gaussian mixture distribution in (3.13) with $d=1000$, $p=0.3$, and $r=[1,1,\ldots,1]\in\mathbb{R}^{d}$. The noise $z$ is sampled from $\mathsf{N}(0,\mathrm{I})$. We compare the linear schedule $\beta_{t}=t$ and the approximate min-avg-Lip2 schedule (3.14); in both cases, $\alpha_{t}=\sqrt{1-\beta_{t}^{2}}$. We use the explicit formula for the drift given in Example 3.12 and run only $2$, 33, or $4$ steps of RK4 to integrate the ODE with $10^{4}$ independent noise samples. For the obtained samples, we use PCA to obtain a 1D projection and fit a 1D bimodal Gaussian mixture model to estimate the weights of the two modes.
Truth Linear schedule Approx min-avg-Lip2 schedule
22 RK4 steps $0.3$ $0.00$ $0.42$
33 RK4 steps $0.3$ $0.03$ $0.26$
44 RK4 steps $0.3$ $0.09$ $0.27$
Table 1. True and estimated weights of one mode recovered from the samples (values reported to 2 decimal places). We obtain two weights since we fit a bimodal GMM, and we always report the smaller weight.
In Table 1, we compare results using the linear schedule and the approximate min-avg-Lip2 schedule. The latter clearly achieves better accuracy, while the former is prone to missing modes.
4.3. Invariant distributions of stochastic Allen-Cahn
We consider an infinite-dimensional probability measure defined over continuous functions on the unit interval $[0,1]$, formally proportional to
(4.1) $\exp\left(-\int_{0}^{1}\frac{1}{2}(\partial_{x}u(x))^{2}+V(u(x)){\rm d}x\right),$
where $V(u)=(1-u^{2})^{2}$ is a double-well potential. This is the stationary distribution of the stochastic Allen-Cahn equation
(4.2) $\partial_{t}u=\partial_{xx}u-V^{\prime}(u)+\sqrt{2}\,\eta\,$,
with natural boundary conditions and space-time white noise $\eta$. The distribution is bimodal, with realizations typically exhibiting rough, approximately constant profiles near $u=\pm 1$. We discretize using finite differences on $N$ equidistributed points, yielding an $N$-dimensional distribution. In the interpolant, we sample $x_{1}$ from this distribution using ensemble MCMC algorithms [7]; $z$ is chosen as white noise. We train an ODE generative model and compare energy spectra between true and generated distributions based on different interpolation schedules. The designed schedule is obtained by simply considering the covariance of the Gaussian measure part $\exp(-\int_{0}^{1}\frac{1}{2}(\partial_{x}u(x))^{2}\,{\rm d}x)$ and applying the optimal avg-Lip2 from the Gaussian case (3.11).
Figure 4. Energy spectra of invariant distributions of stochastic Allen-Cahn: comparison between truth, generated via designed schedules or standard linear schedules, with $10,20$ or $40$ RK4 steps. The three figures correspond to different resolutions. Left: $32$; middle: $64$; right: $128$.
Figure 4 demonstrates that for this mildly non-Gaussian behaved distribution, the designed schedule achieves superior accuracy that remains robust across resolutions, unlike linear schedules.
4.4. Invariant distributions of stochastic Navier-Stokes
Finally, we consider invariant distributions of stochastically forced Navier-Stokes equations on the torus $\mathbb{T}^{2}=[0,2\pi]^{2}$. Using vorticity formulation:
(4.3) $\mathrm{d}\omega+v\cdot\nabla\omega\,\mathrm{d}t=\nu\Delta\omega\,\mathrm{d}t-\alpha\omega\,\mathrm{d}t+\varepsilon\,\mathrm{d}\eta\,$,
where $v=\nabla^{\perp}\psi=(-\partial_{y}\psi,\partial_{x}\psi)$ is the velocity field from stream function $\psi$ satisfying $\Delta\psi=\omega$. We use parameters $\nu=10^{-3}$, $\alpha=0.1$, $\varepsilon=1$, and white-in-time forcing $\mathrm{d}\eta$ on finite Fourier modes, following [8]. The system is ergodic with a unique invariant measure [18].
We generate data for $x_{1}$ by long-time simulation on a fine grid and use white noise for $z$ in the interpolant. For the designed schedule, we observe that at resolution $128\times 128$, the enstrophy spectrum shows $\sim 10^{-4}$ energy at frequency $k=2^{6}$. We apply the schedule from (3.11) with $\lambda^{*}=10^{-5}$. Figure 5 demonstrates that with 10 RK4 steps, the designed schedule produces superior samples with more accurate spectra. Despite the highly non-Gaussian nature of this distribution, schedules optimized for the Gaussian case can still be used to improve fine-scale accuracy.
Figure 5. Left: generated $128\times 128$ sample using linear schedule and $10$ steps of RK4; middle: generated $128\times 128$ sample using designed schedule and $10$ steps of RK4; enstrophy spectra of samples using different schedules.
5. Conclusions
In this paper, we studied the design of interpolation schedules in flow and diffusion-based generative models within the stochastic interpolants framework. We revealed a fundamental paradox: while all scalar interpolation schedules achieve identical statistical efficiency under KL divergence in path space after optimal diffusion tuning, their numerical efficiency can differ dramatically. This statistical equivalence result implies that scalar schedule optimization is inherently limited, and future breakthroughs likely require exploring matrix-valued or nonlinear schedules that could break this equivalence barrier.
To exploit the numerical differences among statistically equivalent scalar schedules, we proposed optimizing averaged squared Lipschitzness of the drift field—a criterion that favors schedules requiring fewer integration steps, contrasting with kinetic energy minimization in optimal transport approaches. Our analytical results demonstrate exponential improvements in Lipschitz constants for Gaussian distributions and reduced mode collapse for mixtures. These insights, derived from simple analytical cases, successfully transfer to complex high-dimensional invariant distributions of stochastic Allen-Cahn and Navier-Stokes equations, achieving robust performance across resolutions.
Our scalar schedule optimization demonstrates meaningful practical improvements across diverse applications, from Gaussian distributions to complex stochastic PDEs. However, the statistical equivalence of scalar schedules points toward matrix-valued interpolation schedules as a natural next step that could unlock significantly greater performance gains. In future work, we plan to explore matrix-valued schedules that can adapt individually to different eigenvalue scales, building on the theoretical foundation established here. Additionally, incorporating physics-informed nonlinear schedules and exploring alternative statistical criteria represent promising avenues for further advancing generative model efficiency.
Acknowledgments YC acknowledges support from UCLA start-up funds. EVE is supported by the National Science Foundation under Awards DMR1420073, DMS-2012510, and DMS-2134216, and by a Vannevar Bush Faculty Fellowship.
Appendix A Sketch of Derivations for Stochastic Interpolants
Sketch of derivation for Proposition 2.2.
For any smooth test function $\phi:\mathbb{R}^{d}\to\mathbb{R}$,
(A.1) ${\rm d}\phi(I_{t})=\dot{I}_{t}\cdot\nabla\phi(I_{t}){\rm d}t\,.$
We denote by $\mu(t,{\rm d}x)$ the measure of $I_{t}$. Then,
(A.2) $\int_{\mathbb{R}^{d}}\phi(x)\mu(t,{\rm d}x)=\mathbb{E}[\phi(I_{t})]=\mathbb{E}[\phi(I_{0})]+\int_{0}^{t}\mathbb{E}[\dot{I}_{s}\cdot\nabla\phi(I_{s})]{\rm d}s\,.$
Using the definition of conditional expectation, we have the identity
(A.3) $\mathbb{E}[\dot{I}_{s}\cdot\nabla\phi(I_{s})]=\mathbb{E}[\mathbb{E}[\dot{I}_{s}|I_{s}]\cdot\nabla\phi(I_{s})]=\int_{\mathbb{R}^{d}}\mathbb{E}[\dot{I}_{s}|I_{s}=x]\cdot\nabla\phi(x)\mu(s,{\rm d}x)\,.$
Combining the above two equations lead to
(A.4) $\int_{\mathbb{R}^{d}}\phi(x)\mu(t,{\rm d}x)=\int_{\mathbb{R}^{d}}\phi(x)\mu(0,{\rm d}x)+\int_{0}^{t}\int_{\mathbb{R}^{d}}\mathbb{E}[\dot{I}_{s}|I_{s}=x]\cdot\nabla\phi(x)\mu(s,{\rm d}x){\rm d}s\,$,
which implies $\mu(t,\cdot)$ is the weak solution to the transport equation corresponding to the ODE ${\rm d}X_{t}=b_{t}(X_{t}){\rm d}t$ with $b_{t}(x)=\mathbb{E}[\dot{I}_{t}|I_{t}=x]$. ∎
Sketch of derivation for Proposition 2.3.
Assume the density of $I_{t}$ exists and denote it by $\rho_{t}$. By Proposition 2.2, $\rho_{t}$ satisfies the transport equation
$\partial_{t}\rho_{t}+\nabla\cdot(\rho_{t}b_{t})=0\,.$
Using the fact that $\nabla\cdot(\rho\nabla\log\rho)=\Delta\rho$, we can rewrite the equation as
$\partial_{t}\rho_{t}+\nabla\cdot(\rho_{t}(b_{t}+\epsilon_{t}\nabla\log\rho_{t}))=\epsilon_{t}\Delta\rho_{t}\,$,
which is exactly the Fokker-Planck equation corresponding to the SDE
${\rm d}X_{t}=\left(b_{t}(X_{t})+\epsilon_{t}\nabla\log\rho_{t}(X_{t})\right){\rm d}t+\sqrt{2\epsilon_{t}}{\rm d}W_{t}\,.$
Sketch of derivation for (2.2).
The second equation in (2.2) follows directly from the first one. Here we derive the first one. Let us denote the density of $\beta_{t}x_{1}$ by $q_{t}$. Then $I_{t}$ is a Gaussian noisy version of $\beta_{t}x_{1}$, implying that
$\rho_{t}(x)\propto\int_{\mathbb{R}^{d}}\rho_{t}(y)\exp(-\frac{\|x-y\|_{2}^{2}}{2\alpha_{t}^{2}}){\rm d}y\,.$
Taking gradient yields the formula
$\nabla\log\rho_{t}(x)=\frac{1}{\int_{\mathbb{R}^{d}}\rho_{t}(y)\exp(-\frac{\|x-y\|_{2}^{2}}{2\alpha_{t}^{2}}){\rm d}y}\int_{\mathbb{R}^{d}}(-\frac{x-y}{\alpha_{t}^{2}})\rho_{t}(y)\exp(-\frac{\|x-y\|_{2}^{2}}{2\alpha_{t}^{2}}){\rm d}y\,.$
On the other hand, by the Bayes rule, we know that
$\frac{1}{\int_{\mathbb{R}^{d}}\rho_{t}(y)\exp(-\frac{\|x-y\|_{2}^{2}}{2\alpha_{t}^{2}}){\rm d}y}\rho_{t}(y)\exp(-\frac{\|x-y\|_{2}^{2}}{2\alpha_{t}^{2}})$
is the density of the conditional distribution $\beta_{t}x_{1}|\alpha_{t}z+\beta_{t}x_{1}=x$. Therefore,
$\nabla\log\rho_{t}(x)=\mathbb{E}[-\frac{x-\beta_{t}x_{1}}{\alpha_{t}^{2}}|I_{t}=x]=-\mathbb{E}[\frac{z}{\alpha_{t}}|I_{t}=x]\,.$
This leads to the first formula in (2.2). ∎
Appendix B Discussion on SDEs with Singular Drift
In Section 2.3, the optimal diffusion coefficient is
$\epsilon_{t}=\alpha_{t}^{2}(\frac{\dot{\beta}_{t}}{\beta_{t}}-\frac{\dot{\alpha}_{t}}{\alpha_{t}})\,.$
With this choice and using the identities in (2.2), we obtain the following SDE
${\rm d}X_{t}=(2b_{t}(X_{t})-\frac{\dot{\beta}_{t}}{\beta_{t}}X_{t}){\rm d}t+\sqrt{2\epsilon_{t}}{\rm d}W_{t}\,.$
For example, we take $\beta_{t}=t,\alpha_{t}=1-t$, which yields
${\rm d}X_{t}=(2b_{t}(X_{t})-\frac{1}{t}X_{t}){\rm d}t+\sqrt{2\frac{1-t}{t}}{\rm d}W_{t}\,.$
The diffusion coefficient is singular and appears worrisome. However, note that
${\rm d}(tX_{t})=2tb_{t}(X_{t}){\rm d}t+\sqrt{2t(1-t)}{\rm d}W_{t}\,$,
which implies that
$X_{t}=\frac{1}{t}\int_{0}^{t}2sb_{s}(X_{s}){\rm d}s+\frac{1}{t}\int_{0}^{t}\sqrt{2s(1-s)}{\rm d}W_{s}\,.$
The last term is well defined as
$\frac{1}{t^{2}}\int_{0}^{t}2s(1-s){\rm d}s=1-\frac{2}{3}t$
is non-singular as $t\to 0$. Therefore, the above stochastic integral equation is well defined. One can use Picard’s iteration to prove the existence of a solution rigorously.
Appendix C Technical Details for Optimizing Averaged Squared Lipschitzness
C.1. Optimal transport drift in the 1D Gaussian case
We provide a sketch of proof for claims made in Remark 3.4. In the Gaussian setting, optimal transport theory implies that the optimal transport map satisfies $Tx=C_{0}^{-\frac{1}{2}}(C_{0}^{\frac{1}{2}}MC_{0}^{\frac{1}{2}})^{\frac{1}{2}}C_{0}^{-\frac{1}{2}}x=\sqrt{\frac{M}{C_{0}}}x$ in 1D. Therefore, the variance at time $t$ in the optimal transport path satisfies
$C_{t}=((1-t)I+tT)C_{0}((1-t)I+tT)^{T}\,.$
Differentiation over $t$ leads to
$\dot{C}_{t}=(T-I)C_{0}((1-t)I+tT)^{T}+((1-t)I+tT)C_{0}(T-I)^{,}.$
On the other hand, let $b_{t}(x)=A_{t}x$, then using the ODE $\dot{x}_{t}=A_{t}x_{t}$ and differentiating $C_{t}=\mathbb{E}[x_{t}x_{t}^{T}]$ leads to the equation $\dot{C}_{t}=A_{t}C_{t}+C_{t}A_{t}^{T}$. Comparing the above two formulas for $\dot{C}_{t}$ implies
$A_{t}=(T-I)((1-t)I+tT)^{-1}\,.$
For 1D, we obtain the formula
$b_{t}(x)=\frac{\sqrt{M}-\sqrt{C_{0}}}{(1-t)\sqrt{C_{0}}+t\sqrt{M}}x\,.$
In particular, we take $C_{0}=1$ to get
$b_{t}(x)=\frac{\sqrt{M}-1}{1-t+t\sqrt{M}}x\,.$
C.2. Formula for Gaussian mixtures
We provide exact formula for the Gaussian mixture model (GMM).
Proposition C.1.
Let the target density be a GMM with $J\in\mathbb{N}$ modes
(C.1) $\rho^{\star}(x)=\sum_{j=1}^{J}p_{j}{\sf N}(x;m_{j},C_{j})$
where $p_{j}\geq 0$ with $\sum_{j=1}^{J}p_{j}=1$, $m_{j}\in\mathbb{R}^{d}$, and $C_{j}=C_{j}^{T}\in\mathbb{R}^{d}\times\mathbb{R}^{d}$ positive-definite. Then
(C.2) $b_{t}(x)$ $=\dot{\beta}_{t}\frac{\sum_{j=1}^{J}p_{j}m_{j}{\sf N}(x;\overline{m}_{j}(t),\overline{C}_{j}(t))}{\sum_{j=1}^{J}p_{j}{\sf N}(x;\overline{m}_{j}(t),\overline{C}_{j}(t))}$
$+\frac{\sum_{j=1}^{J}p_{j}(\beta_{t}\dot{\beta}_{t}C_{j}+\alpha_{t}\dot{\alpha}_{t}I)\overline{C}_{j}^{-1}(t)(x-\overline{m}_{j}(t)){\sf N}(x;\overline{m}_{j}(t),\overline{C}_{j}(t))}{\sum_{j=1}^{J}p_{j}{\sf N}(x;\overline{m}_{j}(t),\overline{C}_{j}(t))}$
where
(C.3) $\overline{m}_{j}(t)=\beta_{t}m_{j},\qquad\overline{C}_{j}(t)=\beta^{2}_{t}C_{j}+\alpha^{2}_{t}\mathrm{I}\,.$
Proof.
By definition
(C.4) $b_{t}(x)$ $=\mathbb{E}[\dot{\beta}_{t}x_{1}+\dot{\alpha}_{t}z|I_{t}=x]$
$=\mathbb{E}[\dot{\beta}_{t}\beta_{t}^{-1}(x-\alpha_{t}z)+\dot{\alpha}_{t}z|I_{t}=x]$
$=\dot{\beta}_{t}\beta^{-1}_{t}x+\alpha_{t}(\alpha_{t}\dot{\beta}_{t}\beta_{t}^{-1}-\dot{\alpha}_{t})\nabla\log\rho_{t}(x)\,.$
where we used the fact $\nabla\log\rho_{t}(x)=-\alpha_{t}^{-1}\mathbb{E}[z|I_{t}=x]$. For the GMM,
(C.5) $\rho_{t}(x)=\sum_{j=1}^{J}p_{j}{\sf N}(x;\overline{m}_{j}(t),\overline{C}_{j}(t))\,$,
so that
(C.6) $\nabla\log\rho_{t}(x)=-\frac{\sum_{j=1}^{J}p_{j}\overline{C}_{j}^{-1}(t)(x-\overline{m}_{j}(t)){\sf N}(x;\overline{m}_{j}(t),\overline{C}_{j}(t))}{\sum_{j=1}^{J}p_{j}{\sf N}(x;\overline{m}_{j}(t),\overline{C}_{j}(t))}\,.$
Inserting this expression in (C.4) we obtain
(C.7) $\frac{\dot{\beta}_{t}}{\beta_{t}}x+\alpha_{t}^{2}\frac{\dot{\beta}_{t}}{\beta_{t}}\nabla\log\rho_{t}(x)$
$=$ $\frac{\dot{\beta}_{t}}{\beta_{t}}\left(x-\frac{\sum_{j=1}^{J}p_{j}(I-\beta^{2}_{t}C_{j}\overline{C}_{j}^{-1}(t))(x-\overline{m}_{j}(t)){\sf N}(x;\overline{m}_{j}(t),\overline{C}_{j}(t))}{\sum_{j=1}^{J}p_{j}{\sf N}(x;\overline{m}_{j}(t),\overline{C}_{j}(t))}\right)$
$=$ $\frac{\dot{\beta}_{t}}{\beta_{t}}\left(\frac{\sum_{j=1}^{J}p_{j}\big{(}\beta_{t}m_{j}+\beta^{2}_{t}C_{j}\overline{C}_{j}^{-1}(x-\overline{m}_{j})\big{)}{\sf N}(x;\overline{m}_{j}(t),\overline{C}_{j}(t))}{\sum_{j=1}^{J}p_{j}{\sf N}(x;\overline{m}_{j}(t),\overline{C}_{j}(t))}\right)$
$=$ $\dot{\beta}_{t}\frac{\sum_{j=1}^{J}p_{j}m_{j}{\sf N}(x;\overline{m}_{j}(t),\overline{C}_{j}(t))}{\sum_{j=1}^{J}p_{j}{\sf N}(x;\overline{m}_{j}(t),\overline{C}_{j}(t))}+\frac{\sum_{j=1}^{J}p_{j}\beta_{t}\dot{\beta}_{t}C_{j}\overline{C}_{j}^{-1}(x-\bar{m}_{j}){\sf N}(x;\overline{m}_{j}(t),\overline{C}_{j}(t))}{\sum_{j=1}^{J}p_{j}{\sf N}(x;\overline{m}_{j}(t),\overline{C}_{j}(t))}\,$,
where in the first and second identities, we used the fact that $\alpha^{2}_{t}\overline{C}_{j}^{-1}(t)=\mathrm{I}-\beta^{2}_{t}C_{j}\overline{C}_{j}^{-1}(t)$.
Now, using $b_{t}(x)=\dot{\beta}_{t}\beta_{t}^{-1}x+\alpha_{t}^{2}(\dot{\beta}_{t}\beta_{t}^{-1}-\dot{\alpha}_{t})\nabla\log\rho_{t}(x)$, we get the final formula. ∎
Remark C.2.
This form of the formula holds generally when $z$ is not of unit covariance. Let $z\sim\mathsf{N}(0,C_{0})$, then we have
(C.8) $b_{t}(x)$ $=\dot{\beta}_{t}\frac{\sum_{j=1}^{J}p_{j}m_{j}{\sf N}(x;\overline{m}_{j}(t),\overline{C}_{j}(t))}{\sum_{j=1}^{J}p_{j}{\sf N}(x;\overline{m}_{j}(t),\overline{C}_{j}(t))}$
$+\frac{\sum_{j=1}^{J}p_{j}(\beta_{t}\dot{\beta}_{t}C_{j}+\alpha_{t}\dot{\alpha}_{t}C_{0})\overline{C}_{j}^{-1}(t)(x-\overline{m}_{j}(t)){\sf N}(x;\overline{m}_{j}(t),\overline{C}_{j}(t))}{\sum_{j=1}^{J}p_{j}{\sf N}(x;\overline{m}_{j}(t),\overline{C}_{j}(t))}$
where
(C.9) $\overline{m}_{j}(t)=\beta_{t}m_{j},\qquad\overline{C}_{j}(t)=\beta^{2}_{t}C_{j}+\alpha^{2}_{t}C_{0}\,.$
When there is only one mode, we get
$b_{t}(x)=\dot{\beta}_{t}m_{1}+(\alpha_{t}\dot{\alpha}_{t}C_{0}+\beta_{t}\dot{\beta}_{t}M)(\alpha_{t}^{2}C_{0}+\beta^{2}_{t}M)^{-1}(x-\beta_{t}m_{1})\,$,
which matches the formula in the Gaussian setting before ($m_{1}=0$).  $\lozenge$
Remark C.3.
Consider the 1D bimodal case
$\mu^{*}(x)=p\mathsf{N}(x;M,1)+(1-p)\mathsf{N}(x;-M,1)\,.$
For general $\alpha_{t},\beta_{t}$, using the formula in Proposition C.1, we have
(C.10) $b_{t}(x)$ $=\dot{\beta}_{t}\frac{pM{\sf N}(x;\beta_{t}M,\beta_{t}^{2}+\alpha_{t}^{2})-(1-p)M{\sf N}(x;-\beta_{t}M,\beta_{t}^{2}+\alpha_{t}^{2})}{p{\sf N}(x;\beta_{t}M,\beta_{t}^{2}+\alpha_{t}^{2})+(1-p){\sf N}(x;-\beta_{t}M,\beta_{t}^{2}+\alpha_{t}^{2})}$
$+(\beta_{t}\dot{\beta}_{t}+\alpha_{t}\dot{\alpha}_{t})(\beta_{t}^{2}+\alpha_{t}^{2})^{-1}\frac{p(x-\beta_{t}M){\sf N}(x;\beta_{t}M,\beta_{t}^{2}+\alpha_{t}^{2})+(1-p)(x+\beta_{t}M){\sf N}(x;\beta_{t}M,\beta_{t}^{2}+\alpha_{t}^{2})}{p{\sf N}(x;\beta_{t}M,\beta_{t}^{2}+\alpha_{t}^{2})+(1-p){\sf N}(x;-\beta_{t}M,\beta_{t}^{2}+\alpha_{t}^{2})}\,.$
Taking $\alpha_{t}=\sqrt{1-\beta_{t}^{2}}$ leads to a simplified formula
$b_{t}(x)$ $=\dot{\beta}_{t}\frac{pM{\sf N}(x;\beta_{t}M,\beta_{t}^{2}+\alpha_{t}^{2})-(1-p)M{\sf N}(x;-\beta_{t}M,\beta_{t}^{2}+\alpha_{t}^{2})}{p{\sf N}(x;\beta_{t}M,\beta_{t}^{2}+\alpha_{t}^{2})+(1-p){\sf N}(x;-\beta_{t}M,\beta_{t}^{2}+\alpha_{t}^{2})}$
$=\dot{\beta}_{t}M\frac{p\exp(2\beta_{t}Mx)-(1-p)}{p\exp(2\beta_{t}Mx)+(1-p)}=\dot{\beta}_{t}M\mathrm{tanh}(h+\beta_{t}Mx)$
where $h$ satisfies $\frac{p}{1-p}=\exp(2h)$ or $p=\frac{\exp(h)}{\exp(h)+\exp(-h)}$.
Moreover, for the $d$ dimensional bimodal Gaussian mixture
$\mu^{*}(x)=p\mathsf{N}(x;r,\mathrm{I})+(1-p)\mathsf{N}(x;-r,\mathrm{I})\,$,
a similar calculation implies $b_{t}(x)=\dot{\beta}_{t}r\mathrm{tanh}(h+\beta_{t}\langle r,x\rangle)$.  $\lozenge$
C.3. Optimizing avg-Lip2 for 1D Gaussian mixtures
Proof for Proposition 3.6.
Using the formula in (3.7), we have $\nabla b_{t}(x)=M^{2}\dot{\beta}_{t}\beta_{t}\operatorname{sech}^{2}(h+\beta_{t}Mx)$ and
(C.11) $A_{2}=\int_{0}^{1}\mathbb{E}[\|\nabla b_{t}(I_{t})\|_{2}^{2}]\,{\rm d}t=M^{4}\int_{0}^{1}\mathbb{E}[\dot{\beta}_{t}^{2}\,\beta_{t}^{2}\,\operatorname{sech}^{4}(h+\beta_{t}\,MI_{t})]{\rm d}t\,.$
We denote $G(u)=\mathbb{E}[\operatorname{sech}^{4}(h+uM(\sqrt{1-u^{2}}z+ux_{1}))]$, so $A_{2}=M^{4}\int_{0}^{1}\dot{\beta}_{t}^{2}\beta_{t}^{2}G(\beta_{t}){\rm d}t$. The Euler-Lagrange equation satisfies
$\frac{{\rm d}}{{\rm d}t}\frac{\partial}{\partial\dot{\beta}_{t}}(\dot{\beta}_{t}^{2}\beta_{t}^{2}G(\beta_{t}))=\frac{\partial}{\partial\beta_{t}}(\dot{\beta}_{t}^{2}\beta_{t}^{2}G(\beta_{t}))\,.$
Using the Beltrami Identity, the equation leads to
$\dot{\beta}_{t}^{2}\beta_{t}^{2}G(\beta_{t})-\dot{\beta}_{t}\frac{\partial}{\partial\dot{\beta}_{t}}(\dot{\beta}_{t}^{2}\beta_{t}^{2}G(\beta_{t}))=\mathrm{const}\,$,
which implies $\dot{\beta}_{t}^{2}\beta_{t}^{2}G(\beta_{t})=\mathrm{const}$ and thus $\dot{\beta}_{t}\beta_{t}(G(\beta_{t}))^{1/2}=\mathrm{const}$. Integrating both sides leads to the solution
$t=\frac{\int_{0}^{\beta_{t}}u(G(u))^{1/2}{\rm d}u}{\int_{0}^{1}u(G(u))^{1/2}{\rm d}u}\,.$
Now, we derive the ODE that $\beta_{t}$ satisfies. To do so, we need to write out the integral over space explicitly. The density of $I_{t}$ satisfies
$\rho_{t}(x)=p\mathsf{N}(x;\beta_{t}M,1)+(1-p)\mathsf{N}(x;-\beta_{t}M,1)\ =\frac{1}{\sqrt{2\pi}}\exp(-\frac{x^{2}+\beta_{t}^{2}M^{2}}{2})\frac{\operatorname{cosh}(h+\beta_{t}Mx)}{\operatorname{cosh}(h)}\,.$
Let us denote $\rho_{t}(x)=\rho(\beta_{t},x)$ in this proof, which allows us to write
(C.12) $A_{2}=M^{4}\int_{0}^{1}\int_{\mathbb{R}}L(\dot{\beta}_{t},\beta_{t},x)\rho(\beta_{t},x){\rm d}x{\rm d}t\,$,
where $L(\dot{\beta}_{t},\beta_{t},x)=\dot{\beta}_{t}^{2}\,\beta_{t}^{2}\,\operatorname{sech}^{4}(h+\beta_{t}\,Mx)$. The Euler-Lagrange equation for this problem has the form
$\int_{\mathbb{R}}\left(\frac{\rm d}{{\rm d}t}\frac{\partial}{\partial\dot{\beta}_{t}}(L(\dot{\beta}_{t},\beta_{t},x)\rho(\beta_{t},x))\right){\rm d}x=\int_{\mathbb{R}}\left(\frac{\partial}{\partial\beta_{t}}(L(\dot{\beta}_{t},\beta_{t},x)\rho(\beta_{t},x))\right){\rm d}x\,.$
We organize the equation according to $\rho$, which leads to
(C.13) $\int_{\mathbb{R}}(\frac{\partial}{\partial\beta_{t}}L-\frac{\rm d}{{\rm d}t}\frac{\partial}{\partial\dot{\beta}_{t}}L)\rho{\rm d}x=\int_{\mathbb{R}}(\frac{\partial}{\partial\dot{\beta}_{t}}L\frac{\rm d}{{\rm d}t}\rho-L\frac{\partial}{\partial\beta_{t}}\rho){\rm d}x\,$,
where we omit the arguments for simplicity of notation.
We have
$\frac{\partial}{\partial\beta_{t}}L=2\dot{\beta}_{t}^{2}\beta_{t}\operatorname{sech}^{4}(h+\beta_{t}Mx)-4Mx\dot{\beta}_{t}^{2}\beta_{t}^{2}\operatorname{sech}^{4}(h+\beta_{t}Mx)\operatorname{tanh}(h+\beta_{t}Mx)$
$\frac{\partial}{\partial\dot{\beta}_{t}}L=2\dot{\beta}_{t}\beta_{t}^{2}\operatorname{sech}^{4}(h+\beta_{t}Mx)$
$\frac{\rm d}{{\rm d}t}\frac{\partial}{\partial\dot{\beta}_{t}}L=(2\ddot{\beta}_{t}\beta_{t}^{2}+4\dot{\beta}_{t}^{2}\beta_{t})\operatorname{sech}^{4}(h+\beta_{t}Mx)-8Mx\dot{\beta}_{t}^{2}\beta_{t}^{2}\operatorname{sech}^{4}(h+\beta_{t}Mx)\operatorname{tanh}(h+\beta_{t}Mx)$
$\frac{\rm d}{{\rm d}t}{\rho}=\dot{\beta}_{t}\frac{\partial}{\partial\beta_{t}}\rho=\dot{\beta}_{t}(-\beta_{t}M^{2}+Mx\operatorname{tanh}(h+\beta_{t}Mx))\rho$
which shows that the left and right hand sides of (C.13) are
$\textrm{LHS}=\int_{\mathbb{R}}\operatorname{sech}^{4}(h+\beta_{t}Mx))\left(-2\dot{\beta}_{t}^{2}\beta_{t}-2\ddot{\beta}_{t}\beta_{t}^{2}+4Mx\dot{\beta}_{t}^{2}\beta_{t}^{2}\operatorname{tanh}(h+\beta_{t}Mx)\right)\rho{\rm d}x\,$,
RHS $=\int_{\mathbb{R}}((\frac{\partial}{\partial\dot{\beta}_{t}}L)\dot{\beta}_{t}-L)\frac{\partial}{\partial\beta_{t}}\rho{\rm d}x=\int_{\mathbb{R}}\left((2\dot{\beta}_{t}\beta_{t}^{2}\operatorname{sech}^{4}(h+\beta_{t}Mx))\dot{\beta}_{t}-L\right)\frac{\partial}{\partial\beta_{t}}\rho{\rm d}x$
$=\int_{\mathbb{R}}\dot{\beta}_{t}^{2}\beta_{t}^{2}\operatorname{sech}^{4}(h+\beta_{t}Mx))\left(-\beta_{t}M^{2}+Mx\operatorname{tanh}(h+\beta_{t}Mx)\right)\rho{\rm d}x\,.$
Since LHS == RHS, we get
$\mathbb{E}[\left(-2\dot{\beta}_{t}^{2}\beta_{t}-2\ddot{\beta}_{t}\beta_{t}^{2}+\dot{\beta}^{2}_{t}\beta^{3}_{t}M^{2}+3\dot{\beta}_{t}^{2}\beta_{t}^{2}MI_{t}\operatorname{tanh}(h+\beta_{t}MI_{t})\right)\operatorname{sech}^{4}(h+\beta_{t}MI_{t})]=0\,.$
Now, we note the fact that $\mathbb{E}[x\operatorname{tanh}(h+\beta_{t}MI_{t})]=\beta_{t}M$ since
$\mathbb{E}[I_{t}\tanh(h+\beta_{t}MI_{t})]$ $=\frac{1}{\sqrt{2\pi}}\int_{\mathbb{R}}e^{-\frac{x^{2}+\beta_{t}^{2}M^{2}}{2}}\frac{\cosh(h+\beta_{t}Mx)}{\cosh(h)}\tanh(h+\beta_{t}Mx)x{\rm d}x$
$=\frac{1}{\sqrt{2\pi}\cosh(h)}\int_{\mathbb{R}}e^{-\frac{x^{2}+\beta_{t}^{2}M^{2}}{2}}\sinh(h+\beta_{t}Mx)x{\rm d}x$
$=\frac{1}{2\sqrt{2\pi}\cosh(h)}\int_{\mathbb{R}}(e^{h}\,e^{-\frac{(x-\beta_{t}M)^{2}}{2}}-e^{-h}\,e^{-\frac{(x+\beta_{t}M)^{2}}{2}})x{\rm d}x$
$=\frac{1}{\sqrt{2\pi}}\cdot\int_{\mathbb{R}}(\frac{e^{h}}{e^{h}+e^{-h}}\,e^{-\frac{(x-\beta_{t}M)^{2}}{2}}-\frac{e^{-h}}{e^{h}+e^{-h}}\,e^{-\frac{(x+\beta_{t}M)^{2}}{2}})x{\rm d}x$
$=\frac{e^{h}}{e^{h}+e^{-h}}\beta_{t}M+\frac{e^{-h}}{e^{h}+e^{-h}}\beta_{t}M=\beta_{t}M\,.$
Thus, we have
$\mathbb{E}[I_{t}\operatorname{tanh}(h+\beta_{t}MI_{t})\operatorname{sech}^{4}(h+mI_{t})]$
$=$ $\operatorname{Cov}(I_{t}\operatorname{tanh}(h+\beta_{t}MI_{t}),\operatorname{sech}^{4}(h+\beta_{t}MI_{t}))+\mathbb{E}[I_{t}\operatorname{tanh}(h+\beta_{t}MI_{t})]\mathbb{E}[\operatorname{sech}^{4}(h+\beta_{t}MI_{t})]$
$=$ $\operatorname{Cov}(I_{t}\operatorname{tanh}(h+\beta_{t}MI_{t}),\operatorname{sech}^{4}(h+\beta_{t}MI_{t}))+\beta_{t}M\mathbb{E}[\operatorname{sech}^{4}(h+\beta_{t}MI_{t})]\,.$
With these formulas, the Euler-Lagrange equation becomes
$2\dot{\beta}_{t}^{2}\beta_{t}-2\ddot{\beta}_{t}\beta_{t}^{2}+\dot{\beta}_{t}^{2}\beta_{t}^{3}M^{2}(4+3\operatorname{Corr}(I_{t}\operatorname{tanh}(h+\beta_{t}MI_{t}),\operatorname{sech}^{4}(h+\beta_{t}MI_{t})))=0\,.$
If we omit the Corr term, we get the ODE
$\dot{\beta}_{t}^{2}\beta_{t}-\ddot{\beta}_{t}\beta_{t}^{2}+2\dot{\beta}_{t}^{2}\beta_{t}^{2}M^{2}=0\,.$
By setting $f_{t}=\beta^{2}_{t}$, the above ODE becomes $\ddot{f}_{t}=M^{2}\dot{f}_{t}$. Solving this ODE with the correct boundary condition leads to
$\beta_{t}=\frac{1}{M}\sqrt{-\log(-M^{2}t+\frac{M^{2}}{1-e^{-M^{2}}})+\log\frac{M^{2}}{1-e^{-M^{2}}}}\,$,
which can be simplified as $\beta_{t}=\frac{1}{M}\sqrt{-\log(1+(e^{-M^{2}}-1)t)}$.
On the other hand, we note that if we optimize $\int_{0}^{1}\mathbb{E}[\|\nabla b_{t}(I_{t})\|^{2k}_{2}]\,\mathrm{d}t$, we will get
$\dot{\beta}_{t}^{2}\beta_{t}-\ddot{\beta}_{t}\beta_{t}^{2}+\dot{\beta}_{t}^{2}\beta_{t}^{3}2M^{2}\left(1+\frac{8k^{2}-6k+1}{8k^{2}-4k}\operatorname{Corr}(I_{t}\operatorname{tanh}(h+\beta_{t}MI_{t}),\operatorname{sech}^{4k}(h+\beta_{t}MI_{t}))\right)=0\,.$
Omitting the correlation part leads to the same equation. Also, using the argument at the beginning of this proof, we have in such case
$t=\frac{\int_{0}^{\beta_{t}}u(G(u))^{1/{2k}}{\rm d}u}{\int_{0}^{1}u(G(u))^{1/{2k}}{\rm d}u}\,$,
where $G(u)=\mathbb{E}[\operatorname{sech}^{4k}(h+uM(\sqrt{1-u^{2}}z+ux_{1}))]$.

### v2

@@
 We show that while all scalar interpolation schedules achieve identical statistical efficiency under Kullback-Leibler divergence in path space after optimal diffusion coefficient tuning, their numerical efficiency can differ substantially.
-This observation motivates focusing on numerical properties of the resulting drift fields rather than statistical criteria for schedule design.
+This motivates focusing on numerical properties of the resulting drift fields rather than purely statistical criteria for schedule design.
 We propose averaged squared Lipschitzness minimization as a principled criterion for numerical optimization, providing an alternative to kinetic energy minimization used in optimal transport approaches.
 A transfer formula is derived that enables conversion between different schedules at inference time without retraining neural networks.
-For Gaussian distributions, our optimized schedules achieve exponential improvements in Lipschitz constants over standard linear schedules, while for Gaussian mixtures, they reduce mode collapse in few-step sampling.
+For Gaussian distributions, the optimized schedules achieve exponential improvements in Lipschitz constants over standard linear schedules, while for Gaussian mixtures, they reduce mode collapse in few-step sampling.
 We also validate our approach on high-dimensional invariant distributions from stochastic Allen-Cahn equations and Navier-Stokes equations, demonstrating robust performance improvements across resolutions.
@@
 1.1. Context
-Dynamics between probability measures, particularly flows and diffusion processes described by ordinary and stochastic differential equations (ODEs and SDEs), form the foundation of state-of-the-art generative modeling techniques [38, 19, 42].
-These methods generate samples through an iterative refinement process that progressively eliminates noise or corruption at different scales [40, 23].
-In this paper, we study the impact and design of interpolation schedules on the performance of flow and diffusion-based generative models.
-We work within the stochastic interpolant framework [1, 2], which provides a systematic approach for modeling noising processes through sample interpolation and enables principled construction of the corresponding generative processes.
-This framework connects to related concurrent work on flow matching [27] and rectified flows [28], and encompasses diffusion and score-based generative models [38, 42, 19, 41] as specific instances.
+Dynamics between probability measures, particularly flows and diffusion processes described by ordinary and stochastic differential equations (ODEs and SDEs), underpin many modern generative modeling techniques [sohl2015deep, ho2020denoising, song2020score].
+These models generate samples by progressively transforming simple noise into structured data through a sequence of intermediate distributions, implemented as an iterative denoising or refinement process [song2019generative, karras2022elucidating].
+The manner in which these transformations are traversed in time—encoded by interpolation or noise schedules—plays a crucial role in shaping the resulting generative dynamics and their performance.
+In this paper, we study the mathematical design of interpolation schedules for flow and diffusion-based generative models within the stochastic interpolants framework [albergo2023stochastic, albergo2022building].
+This framework provides a unified formulation for noising and denoising processes based on sample interpolation, and enables principled construction of the associated generative dynamics.
+It is closely related to concurrent developments such as flow matching [lipman2022flow] and rectified flows [liu2022flow], and encompasses diffusion and score-based generative models [sohl2015deep, song2020score, ho2020denoising, song2020improved] as special cases.
 1.2. Basics of stochastic interpolants
@@
 For different values of $t$, the interpolant $I_{t}$ can be seen as modeling a corruption of the target at a specific scale.
-The theory of stochastic interpolants [2, 1] shows that one can generate samples from $\mu^{*}$ by solving the following ODE:
+The theory of stochastic interpolants [albergo2022building, albergo2023stochastic] shows that one can generate samples from $\mu^{*}$ by solving the following ODE:
 (1.1) $\mathrm{d}X_{t}=b_{t}(X_{t})\mathrm{d}t,\quad X_{0}\sim\mathsf{N}(0,\mathrm{I})\,$,
@@
 The solution satisfies $\mathrm{Law}(X_{t})=\mathrm{Law}(I_{t})$, and in particular, $X_{1}\sim\mu^{*}$.
-This can also be seen as a consequence of the mimicking theorem [17], also referred to as Markovian projection.
+This can also be seen as a consequence of the mimicking theorem [gyongy1986mimicking], also referred to as Markovian projection.
 Because the drift $b_{t}$ is a conditional expectation, we can define it as the minimizer of the square loss function
@@
   • In Section 2, we prove that under the Kullback-Leibler divergence criterion in path space, different choices of scalar schedules are statistically equivalent when diffusion coefficients are optimized a posteriori.
-This equivalence paradoxically renders statistical considerations insufficient for schedule selection.
+This equivalence fundamentally renders statistical considerations insufficient for schedule selection.
   • In Sections 3.1 and 3.2, we introduce a principled approach to numerical optimization by minimizing the averaged squared Lipschitzness of the drift $b_{t}$ at inference time.
@@
 1.4. Related work
-Since the introduction of flow and diffusion models, numerous studies have examined the design principles and parameter space of these models (see a review in [48]).
-These investigations encompass the choice of noise, noising processes, time reversal processes, training losses, and diffusion coefficients in the generative process.
-This paper focuses on designing interpolation schedules within the framework of unit-time stochastic interpolants [2, 1], which relates to the noising and denoising schedules in diffusion models.
-Schedule design is of interest from both statistical and numerical perspectives.
-From a statistical standpoint, it was demonstrated in [25] that different noise schedules in diffusion models yield the same variational lower bound.
-Our results suggest that this “statistical equivalence” generalizes to a broader context using the unit-time stochastic interpolants framework, with the Kullback-Leibler divergence in path space serving as the statistical estimation criterion (see discussion in Remark 2.6).
-From a numerical perspective, existing works have derived insights primarily through empirical studies on machine learning datasets to tune the noise schedules for efficient sampling performance [35, 22, 30, 41, 23]; see also [36, 47, 32, 5] for learning improved schedules with additional training and an analysis [45] considering score errors.
-In this work, we propose a principled way for numerical design by optimizing the Lipschitzness of the drift field at inference time.
-Related mathematical work focused on the Lipschitz regularity of flows and flow maps includes [9, 44].
-See also [3] for a mathematical investigation of the numerical impact of schedules on identifying modes in high dimensions.
-It has been advocated to learn the optimal transport path [28], which is straight and therefore offers better numerical performance; see also generative models built using entropy-regularized optimal transport, namely Schrödinger bridges [11, 37].
-Nevertheless, the optimal transport path may lead to irregular drift fields [44] that are not ideal for numerical integration (see also an example in Remark 3.4), an issue our proposed criterion of optimizing the Lipschitzness aims to address.
-Moreover, there has been a line of work on improving numerical performance with high order, exponential, or parallel integrators, e.g., [13, 29, 50, 26, 6, 46, 10, 43], and multiscale and cascading approach [49, 12, 21, 33, 20, 16, 31], which can be combined with the design of schedules to accelerate sampling.
-We also note another line of work focused on consistency models and learning flow maps (e.g., [39, 24, 34, 14, 4]), to achieve few steps sampling and thus improve numerical efficiency.
+Since the introduction of flow and diffusion-based generative models, a large body of work has explored their design principles and parameter choices; see the survey [yang2023diffusion].
+These studies address, among other aspects, the choice of noise distributions, noising and denoising processes, time-reversal dynamics, training objectives, and diffusion coefficients.
+The present work focuses on the design of interpolation schedules within the unit-time stochastic interpolants framework [albergo2022building, albergo2023stochastic], which relates to concurrent developments in flow matching [lipman2022flow] and rectified flows [liu2022flow], and encompasses diffusion and score-based generative models [sohl2015deep, song2020score, ho2020denoising, song2020improved] as special cases.
+Within this framework, interpolation schedules play a role analogous to noise schedules in diffusion models.
+Schedule design has been studied from both statistical and numerical perspectives.
+From a statistical standpoint, [kingma2021variational] showed that different noise schedules in diffusion models yield the same variational lower bound.
+Our results indicate that this notion of statistical equivalence extends to a broader setting within the unit-time stochastic interpolants framework when the Kullback–Leibler divergence in path space is used as the estimation criterion; see Remark 2.6.
+This observation suggests that statistical considerations alone are insufficient to distinguish among scalar interpolation schedules.
+From a numerical perspective, most existing work has relied on empirical studies—primarily on machine learning benchmarks—to tune noise or time schedules for improved sampling efficiency [san2021noise, jolicoeur2021gotta, nichol2021improved, song2020improved, karras2022elucidating].
+Related approaches learn improved schedules or time parametrizations using additional training [shaul2024bespoke, xue2024accelerating, sabouralign, chentrajectory], while [wang2024evaluating] analyzes the sensitivity of schedule choice to score estimation errors.
+Here we propose a principled approach to numerical schedule design based on optimizing the Lipschitz regularity of the drift field at inference time, without requiring retraining.
+Related mathematical work has investigated Lipschitz regularity, contractivity, and stability properties of flows and flow maps [daniels2025contractivity, tsimpos2025optimal].
+In particular, [tsimpos2025optimal] considers a variational problem for minimizing the maximum Lipschitz constant under reparametrization of a fixed flow, a setting which differs from ours.
+See also [aranguri2025optimizing] for a mathematical study of how schedule design affects mode identification in high-dimensional distributions.
+Another closely related line of work advocates learning optimal transport paths [liu2022flow, pooladian2023multisample], which are straight and therefore appealing from a numerical standpoint; related models based on entropy-regularized optimal transport include Schrödinger bridges [de2021diffusion, shi2023diffusion, pooladian2025plug].
+However, optimal transport paths may give rise to highly irregular drift fields [tsimpos2025optimal], which are not well suited for numerical integration (see also Remark 3.4).
+Moreover, since the estimation of the initial drift $b_{0}$ inevitably contains errors, one-step or few-step generation strategies that rely heavily on straightness may struggle to accurately reproduce fine-scale features.
+Our proposed Lipschitz-based criterion is instead designed to directly target numerical integration efficiency and robustness to discretization.
+Finally, numerical efficiency can also be improved through advances orthogonal to schedule design, including higher-order, exponential, and parallel integrators [dockhorn2022genie, lu2022dpm, zhang2022fast, li2024accelerating, chen2024accelerating, wu2024stochastic, de2025accelerated, tan2025stork], as well as multiscale and cascading approaches [yu2020wavelet, dhariwal2021diffusion, jing2022subspace, saharia2022image, ho2022cascaded, guth2022wavelet, phung2023wavelet].
+In addition, consistency models and approaches based on learning flow maps [song2023consistency, kim2024consistency, salimans2022progressive, frans2024one, boffi2025flow] aim to reduce the number of sampling steps altogether.
+These approaches are complementary to schedule design and can be combined with the methods studied in this work.
+We note that interpolation schedules can also substantially influence training stability and efficiency.
+The present paper focuses primarily on the statistical and numerical efficiency of schedule design.
 2. Statistical Equivalence under Kullback-Leibler in Path Space
@@
 2.1. Stochastic interpolants
-Here we briefly recall the main results of the stochastic interpolant framework [2, 1]. For completeness, we also include a simple sketch of derivations in Appendix A.
+Here we briefly recall the main results of the stochastic interpolant framework [albergo2022building, albergo2023stochastic].
+For completeness, we also include a simple sketch of derivations in Appendix A.
 As in Section 1.2, we denote the target distribution by $\mu^{*}$, and assume that it is supported on $\mathbb{R}^{d}$ and satisfies $\mathbb{E}[\|x_{1}\|_{2}^{2}]<\infty$.
@@
 Let us consider the criterion of the KL divergence between path measures $\mathbb{P}_{X}$ and $\mathbb{P}_{\hat{X}}$ of $X=(X_{t})_{0\leq t\leq 1}$ and $\hat{X}=(\hat{X}_{t})_{0\leq t\leq 1}$, respectively.
-Aaccording to Girsanov’s theorem, this KL divergence has the form
-(2.5) $\mathrm{KL}[\mathbb{P}_{X}\|\mathbb{P}_{\hat{X}}]=\frac{1}{2\epsilon_{t}}\int_{0}^{1}\left(\alpha_{t}^{2}(\frac{\dot{\beta}_{t}}{\beta_{t}}-\frac{\dot{\alpha}_{t}}{\alpha_{t}})+\epsilon_{t}\right)^{2}\|\nabla\log\rho_{t}(x)-\hat{s}_{t}(x)\|^{2}_{2}\rho_{t}(x){\rm d}t\,.$
+According to Girsanov’s theorem, this KL divergence has the form
+(2.5) $\mathrm{KL}[\mathbb{P}_{X}\|\mathbb{P}_{\hat{X}}]=\frac{1}{2\epsilon_{t}}\int_{\mathbb{R}^{d}}\int_{0}^{1}\left(\alpha_{t}^{2}(\frac{\dot{\beta}_{t}}{\beta_{t}}-\frac{\dot{\alpha}_{t}}{\alpha_{t}})+\epsilon_{t}\right)^{2}\|\nabla\log\rho_{t}(x)-\hat{s}_{t}(x)\|^{2}_{2}\rho_{t}(x){\rm d}t{\rm d}x\,.$
 Now, recall the fact that, for any $a$, the minimizer of $\frac{(\epsilon+a)^{2}}{2\epsilon}=\frac{\epsilon}{2}+a+\frac{a^{2}}{2\epsilon}$ is $\epsilon=|a|$, and the minimum is $\max\{0,2a\}$.
@@
 Remark 2.6.
-In [25], it was pointed out that in diffusion models, different noise schedules lead to the same variational lower bound.
+In [kingma2021variational], it was pointed out that in diffusion models, different noise schedules lead to the same variational lower bound.
 In the continuous setting, this corresponds to the KL divergence in path space.
@@
 $\lozenge$
-Proposition 2.5 shows that the optimal KL accuracy in path space depends solely on the estimation of $\nabla\log q_{r}(x_{1}+rz)$: that is, from the perspective of KL divergence in path space, all linear scalar interpolants with independently coupled endpoints and one endpoint Gaussian are statistically indistinguishable.
+Proposition 2.5 shows that the optimal KL accuracy in path space depends solely on the estimation of $\nabla\log q_{r}(x_{1}+rz)$: that is, from the perspective of KL divergence in path space, all linear scalar interpolants with independently coupled endpoints and one endpoint being Gaussian are statistically indistinguishable.
 This indicates that other metrics need to be explored if we want to select models for improved statistical efficiency.
-On the other hand, using matrix-valued instead of scalar schedules may potentially lead to different statistical efficiency, a direction of interest in future work.
+On the other hand, using nonlinear or matrix-valued instead of scalar schedules may potentially lead to different statistical efficiency, a direction of interest in future work.
 3. Numerical Design by Optimizing Averaged Squared Lipschitzness
@@
 In this section, we explore how to choose interpolation schedules that enhance numerical efficiency.
-We focus on ODEs rather than SDEs for simplicity, noting that ODEs typically achieve better empirical performance due to their greater ease of integration [23, 13].
+We focus on ODEs rather than SDEs for simplicity, noting that ODEs typically achieve better empirical performance due to their greater ease of integration [karras2022elucidating, dockhorn2022genie].
 3.1. From one schedule to another
@@
 This also applies to the estimators of the drift functions, so we can tune the schedule at inference time rather than during training.
-Similar statements have appeared in the literature [25, 23].
+Similar statements have appeared in the literature [kingma2021variational, karras2022elucidating].
 We will use this fact in numerical experiments in Section 4.
@@
 Remark 3.7.
-The time-dilated schedule studied in [3] also resolve the mode missing issue:
+The time-dilated schedule studied in [aranguri2025optimizing] also resolve the mode missing issue:
 (3.10) $\beta_{t}=\begin{cases}\displaystyle\frac{2\kappa\,t}{M},&t\in\bigl[0,\tfrac{1}{2}\bigr]\,,\\[4.30554pt] \displaystyle\frac{\kappa}{M}+\Bigl(1-\frac{\kappa}{M}\Bigr)\,(2t-1),&t\in\bigl[\tfrac{1}{2},1\bigr]\,.\end{cases}$
@@
 $\frac{\alpha_{t}\dot{\alpha}_{t}+\beta_{t}\dot{\beta}_{t}\lambda_{M}^{-1}}{\alpha_{t}^{2}+\beta_{t}^{2}\lambda_{M}^{-1}}\preceq\nabla b_{t}(x)\preceq\frac{\alpha_{t}\dot{\alpha}_{t}+\beta_{t}\dot{\beta}_{t}\lambda_{m}^{-1}}{\alpha_{t}^{2}+\beta_{t}^{2}\lambda_{m}^{-1}}\,.$
-This can be proved using the Cramér–Rao and Brascamp–Lieb inequalities; see [15]. Therefore, similar to the Gaussian case, we can choose $\lambda^{\star}=\lambda_{M}^{-1}$. Then, with the schedule
+This can be proved using the Cramér–Rao and Brascamp–Lieb inequalities; see [gao2024gaussian].
+Therefore, similar to the Gaussian case, we can choose $\lambda^{\star}=\lambda_{M}^{-1}$.
+Then, with the schedule
 (3.12) $\alpha_{t}=\sqrt{\frac{\lambda^{\star}-(\lambda^{\star})^{t}}{\lambda^{\star}-1}},\quad\beta_{t}=\sqrt{\frac{(\lambda^{\star})^{t}-1}{\lambda^{\star}-1}}\,$,
@@
 4. Numerical Demonstrations
-In this section, we conduct numerical experiments to demonstrate the improved efficiency of the schedule designed in the previous section.
-We first study high-dimensional Gaussian distributions and Gaussian mixtures, followed by the high-dimensional invariant distributions of the stochastic Allen-Cahn equation and the Navier-Stokes equations, which exhibit slightly and highly non-Gaussian behaviors respectively.
-We use the UNet architecture popularized by [19] to train the drift field for all experiments except Gaussian and mixtures, where we employ explicit formula.
-In all the examples, we integrate the ODE from $t_{\min}=10^{-3}$ to $t_{\max}=1-10^{-3}$ to avoid potential numerical issues at $t=0$ or $1$.
-Code is available at https://github.com/yifanc96/GenerativeDynamics-NumericalDesign.git.
-For accuracy evaluation, we use the energy spectrum (or enstrophy spectrum in the case of Navier-Stokes) of the samples as the criterion.
-The spectrum for a sample $u$ (which is a function that is either 1D or 2D in this paper) is computed using the formula
-$E(k)=\sum_{k\leq|m|_{2}\leq k+1}|\hat{u}(m)|^{2}\,$,
-where $\hat{u}(m)$ are the Fourier coefficients. We average $E(k)$ over sufficiently many samples for each frequency $k$.
+In this section, we conduct numerical experiments to demonstrate the improved numerical efficiency of the interpolation schedules designed in the previous section.
+We consider four classes of target distributions of increasing complexity: high-dimensional Gaussian random fields, high-dimensional Gaussian mixtures, invariant distributions of the stochastic Allen–Cahn equation, and invariant distributions of the stochastic Navier–Stokes equations.
+These examples range from exactly Gaussian to strongly non-Gaussian and serve to illustrate the robustness of the proposed schedule design principles.
+In all experiments, the generative process is formulated as an ODE derived from stochastic interpolants.
+Except for the Gaussian and Gaussian mixture cases, where the drift fields are available in closed form, we parameterize the drift using the UNet architecture introduced in [ho2020denoising].
+Training procedures, network architectures, and optimization hyperparameters follow [chen2024probabilistic] and are kept identical across different interpolation schedules.
+For numerical stability, the ODE is integrated from $t_{\min}=10^{-3}$ to $t_{\max}=1-10^{-3}$ in all experiments.
+Code to reproduce all results is available at https://github.com/yifanc96/GenerativeDynamics-NumericalDesign.git.
+Sampling accuracy is evaluated using spectral diagnostics. For a generated sample $u$ (a one- or two-dimensional function in this paper), we compute the energy spectrum
+$E(k)=\sum_{k\leq|m|_{2}<k+1}|\hat{u}(m)|^{2},$
+where $\hat{u}(m)$ denotes the Fourier coefficients.
+In the Navier–Stokes case with vorticity formulation, this corresponds to the enstrophy spectrum.
+Reported spectra are obtained by averaging over a sufficiently large ensemble of independently generated samples.
 4.1. Gaussians
-We consider the Gaussian random field $\mathsf{N}(0,\sigma^{2}(-\Delta+\tau^{2}I)^{-s})$, where $\Delta$ is the negative Laplacian with homogeneous Dirichlet boundary conditions on $D=[0,1]^{2}$.
-The true data distribution $x_{1}$ is sampled from this distribution with parameters $s=3$, $\tau=1$, and $\sigma^{2}=(4\pi^{2}+\tau^{2})^{s}$.
-The noise $z$ in the interpolant is sampled from white noise (which corresponds to $\sigma=1,s=0$ in the Gaussian random field).
-We discretize the 2D field on a grid with $N$ points in each dimension and construct flow-based generative models.
-The ODE is solved using the fourth-order Runge-Kutta (RK4) scheme.
-Figure 2 shows random fields generated using a standard linear schedule $\beta_{t}=t$ compared to those using our designed schedule (3.11) optimized for avg-Lip2, both employing 20 RK4 steps with $N=128$.
-The designed schedule clearly produces superior samples.
-The right panel of Figure 2 displays both schedules: notably, the designed schedule exhibits rapid initial growth.
-In Figure 3, we compare the energy spectra of the true distribution and generated samples.
-The designed schedule yields a more accurate spectrum, and this accuracy remains robust as resolution increases, unlike standard linear schedules where performance degrades with refinement.
+We first consider a Gaussian random field with distribution $\mathsf{N}(0,\sigma^{2}(-\Delta+\tau^{2}\mathrm{I})^{-s})$ on $D=[0,1]^{2}$, where $\Delta$ denotes the Laplacian with homogeneous Dirichlet boundary conditions.
+We fix $s=3$, $\tau=1$, and $\sigma^{2}=(4\pi^{2}+\tau^{2})^{s}$, so that the field exhibits significant scale separation across Fourier modes.
+Target samples $x_{1}$ are drawn exactly from this distribution.
+The noise variable $z$ in the stochastic interpolant is sampled from spatial white noise, corresponding to the Gaussian random field with $s=0$ and unit variance.
+The field is discretized on a uniform $N\times N$ grid, and the resulting generative ODE is integrated using a fixed-step fourth-order Runge–Kutta (RK4) method.
+We compare the standard linear schedule $\beta_{t}=t$ with the designed schedule (3.11), which minimizes avg-Lip2 for Gaussian targets.
+Figure 2 shows representative samples at resolution $N=128$ generated using 20 RK4 steps.
+The designed schedule produces visibly smoother and more coherent samples.
+The right panel illustrates the schedules themselves, highlighting the rapid initial growth of the designed schedule, which reflects the fast speed required at early times.
+Figure 3 compares the energy spectra of the true distribution and generated samples across resolutions.
+The designed schedule yields significantly more accurate spectra and maintains accuracy as $N$ increases, whereas the linear schedule deteriorates under refinement, reflecting its poorer numerical conditioning.
 Figure 2.
@@
 4.2. Gaussian mixtures
-We consider the $d$-dimensional Gaussian mixture distribution in (3.13) with $d=1000$, $p=0.3$, and $r=[1,1,\ldots,1]\in\mathbb{R}^{d}$.
-The noise $z$ is sampled from $\mathsf{N}(0,\mathrm{I})$.
-We compare the linear schedule $\beta_{t}=t$ and the approximate min-avg-Lip2 schedule (3.14); in both cases, $\alpha_{t}=\sqrt{1-\beta_{t}^{2}}$.
-We use the explicit formula for the drift given in Example 3.12 and run only $2$, 33, or $4$ steps of RK4 to integrate the ODE with $10^{4}$ independent noise samples.
-For the obtained samples, we use PCA to obtain a 1D projection and fit a 1D bimodal Gaussian mixture model to estimate the weights of the two modes.
+We next consider the high-dimensional Gaussian mixture distribution defined in (3.13) with dimension $d=1000$, mixture weight $p=0.3$, and mean vector $r=(1,1,\ldots,1)\in\mathbb{R}^{d}$.
+This distribution is strongly non-Gaussian and exhibits well-separated modes.
+The noise variable $z$ is sampled from $\mathsf{N}(0,\mathrm{I})$.
+We compare the linear schedule $\beta_{t}=t$ with the approximate min-avg-Lip2 schedule (3.14), taking $\alpha_{t}=\sqrt{1-\beta_{t}^{2}}$ in both cases.
+The drift field is given analytically by Example 3.12, allowing us to isolate the numerical effects of the schedule without learning error.
+The ODE is integrated using only 2, 3, or 4 RK4 steps, with $10^{4}$ independent samples generated for each setting.
+To assess mode recovery, we project the generated samples onto one dimension using PCA and fit a one-dimensional bimodal Gaussian mixture model.
+Table 1 reports the estimated smaller mixture weight.
+The linear schedule frequently collapses to a single mode when using few steps, while the approximate min-avg-Lip2 schedule accurately recovers both modes even under extremely coarse discretization.
 Truth Linear schedule Approx min-avg-Lip2 schedule
@@
 We obtain two weights since we fit a bimodal GMM, and we always report the smaller weight.
-In Table 1, we compare results using the linear schedule and the approximate min-avg-Lip2 schedule. The latter clearly achieves better accuracy, while the former is prone to missing modes.
 4.3. Invariant distributions of stochastic Allen-Cahn
-We consider an infinite-dimensional probability measure defined over continuous functions on the unit interval $[0,1]$, formally proportional to
-(4.1) $\exp\left(-\int_{0}^{1}\frac{1}{2}(\partial_{x}u(x))^{2}+V(u(x)){\rm d}x\right),$
-where $V(u)=(1-u^{2})^{2}$ is a double-well potential. This is the stationary distribution of the stochastic Allen-Cahn equation
-(4.2) $\partial_{t}u=\partial_{xx}u-V^{\prime}(u)+\sqrt{2}\,\eta\,$,
-with natural boundary conditions and space-time white noise $\eta$.
-The distribution is bimodal, with realizations typically exhibiting rough, approximately constant profiles near $u=\pm 1$.
-We discretize using finite differences on $N$ equidistributed points, yielding an $N$-dimensional distribution.
-In the interpolant, we sample $x_{1}$ from this distribution using ensemble MCMC algorithms [7]; $z$ is chosen as white noise.
-We train an ODE generative model and compare energy spectra between true and generated distributions based on different interpolation schedules.
-The designed schedule is obtained by simply considering the covariance of the Gaussian measure part $\exp(-\int_{0}^{1}\frac{1}{2}(\partial_{x}u(x))^{2}\,{\rm d}x)$ and applying the optimal avg-Lip2 from the Gaussian case (3.11).
+We consider the invariant distribution of the stochastic Allen–Cahn equation on $[0,1]$, formally given by
+$\exp\!\left(-\int_{0}^{1}\frac{1}{2}(\partial_{x}u(x))^{2}+V(u(x))\,\mathrm{d}x\right),\qquad V(u)=(1-u^{2})^{2}.$
+This distribution is bimodal and moderately non-Gaussian, with samples concentrating near $u=\pm 1$.
+We discretize the equation using finite differences on $N$ equispaced grid points, yielding an $N$-dimensional target distribution.
+Samples $x_{1}$ from the invariant distribution are generated using ensemble MCMC methods [chen2025new], while $z$ is sampled from spatial white noise.
+We train an ODE-based generative model and compare the energy spectra of samples generated using different interpolation schedules.
+The designed schedule is constructed by considering the covariance of the Gaussian reference measure $\exp(-\int_{0}^{1}\tfrac{1}{2}(\partial_{x}u)^{2}\,\mathrm{d}x)$ and applying the Gaussian avg-Lip2-optimal schedule (3.11).
+Figure 4 shows that this schedule yields consistently more accurate energy spectra than the linear schedule and that this improvement is robust across spatial resolutions.
 Figure 4.
@@
 Left: $32$; middle: $64$; right: $128$.
-Figure 4 demonstrates that for this mildly non-Gaussian behaved distribution, the designed schedule achieves superior accuracy that remains robust across resolutions, unlike linear schedules.
 4.4. Invariant distributions of stochastic Navier-Stokes
-Finally, we consider invariant distributions of stochastically forced Navier-Stokes equations on the torus $\mathbb{T}^{2}=[0,2\pi]^{2}$. Using vorticity formulation:
+Finally, we consider invariant distributions of the two-dimensional stochastically forced Navier–Stokes equations on the torus $\mathbb{T}^{2}=[0,2\pi]^{2}$.
+Using the vorticity formulation, the dynamics are given by
 (4.1) $\mathrm{d}\omega+v\cdot\nabla\omega\,\mathrm{d}t=\nu\Delta\omega\,\mathrm{d}t-\alpha\omega\,\mathrm{d}t+\varepsilon\,\mathrm{d}\eta,$
-where $v=\nabla^{\perp}\psi=(-\partial_{y}\psi,\partial_{x}\psi)$ is the velocity field from stream function $\psi$ satisfying $\Delta\psi=\omega$.
-We use parameters $\nu=10^{-3}$, $\alpha=0.1$, $\varepsilon=1$, and white-in-time forcing $\mathrm{d}\eta$ on finite Fourier modes, following [8].
-The system is ergodic with a unique invariant measure [18].
-We generate data for $x_{1}$ by long-time simulation on a fine grid and use white noise for $z$ in the interpolant.
-For the designed schedule, we observe that at resolution $128\times 128$, the enstrophy spectrum shows $\sim 10^{-4}$ energy at frequency $k=2^{6}$.
-We apply the schedule from (3.11) with $\lambda^{*}=10^{-5}$.
-Figure 5 demonstrates that with 10 RK4 steps, the designed schedule produces superior samples with more accurate spectra.
-Despite the highly non-Gaussian nature of this distribution, schedules optimized for the Gaussian case can still be used to improve fine-scale accuracy.
+where $v=\nabla^{\perp}\psi=(-\partial_{y}\psi,\partial_{x}\psi)$ is the incompressible velocity field associated with the stream function $\psi$ satisfying $\Delta\psi=\omega$.
+We fix the parameters to $\nu=10^{-3}$, $\alpha=0.1$, and $\varepsilon=1$.
+The stochastic forcing $\eta$ is white in time and acts on a finite number of low-frequency Fourier modes, following the setup in [chen2024probabilistic].
+Under these choices, the system is ergodic and admits a unique invariant probability measure [hairer2006ergodicity].
+Samples from the invariant distribution are generated via long-time numerical simulation of (4.1) using a pseudo-spectral method with standard de-aliasing.
+After an initial burn-in period sufficient for equilibration, vorticity snapshots are collected at regular time intervals to form an approximately independent dataset of samples.
+In the stochastic interpolant framework, samples $x_{1}$ are drawn from this empirical invariant distribution, while the initial condition $z$ is sampled from spatial white noise.
+We train an ODE-based generative model to transport $z$ to $x_{1}$ over unit time using a learned drift field.
+The drift is parameterized by a UNet architecture, and training follows the same procedure as in previous sections.
+To investigate the effect of interpolation schedules, we compare the standard linear schedule $\beta_{t}=t$ with a Lipschitz-optimized schedule motivated by the Gaussian analysis in Section 3.5.
+Specifically, we apply the schedule in (3.11) with $\lambda^{\star}=10^{-5}$.
+This choice is motivated by the empirical observation that, at resolution $128\times 128$, the enstrophy spectrum of the invariant distribution has magnitude on the order of $10^{-4}$ at frequency $k=2^{6}$, while white noise has magnitude around $10$ there, so take $\lambda^{\star}=10^{-4}/10=10^{-5}$.
+The ODE is solved using a fixed-step fourth-order Runge–Kutta (RK4) method with a small number of uniform time steps.
+Figure 5 shows that with only 10 RK4 steps, the designed schedule produces samples with significantly more accurate enstrophy spectra than the linear schedule.
+Despite the highly non-Gaussian nature of the invariant measure, schedules optimized for the Gaussian case still yield substantial improvements in fine-scale accuracy.
 Figure 5.
@@
 5. Conclusions
-In this paper, we studied the design of interpolation schedules in flow and diffusion-based generative models within the stochastic interpolants framework.
-We revealed a fundamental paradox: while all scalar interpolation schedules achieve identical statistical efficiency under KL divergence in path space after optimal diffusion tuning, their numerical efficiency can differ dramatically.
-This statistical equivalence result implies that scalar schedule optimization is inherently limited, and future breakthroughs likely require exploring matrix-valued or nonlinear schedules that could break this equivalence barrier.
-To exploit the numerical differences among statistically equivalent scalar schedules, we proposed optimizing averaged squared Lipschitzness of the drift field—a criterion that favors schedules requiring fewer integration steps, contrasting with kinetic energy minimization in optimal transport approaches.
-Our analytical results demonstrate exponential improvements in Lipschitz constants for Gaussian distributions and reduced mode collapse for mixtures.
-These insights, derived from simple analytical cases, successfully transfer to complex high-dimensional invariant distributions of stochastic Allen-Cahn and Navier-Stokes equations, achieving robust performance across resolutions.
-Our scalar schedule optimization demonstrates meaningful practical improvements across diverse applications, from Gaussian distributions to complex stochastic PDEs.
-However, the statistical equivalence of scalar schedules points toward matrix-valued interpolation schedules as a natural next step that could unlock significantly greater performance gains.
-In future work, we plan to explore matrix-valued schedules that can adapt individually to different eigenvalue scales, building on the theoretical foundation established here.
-Additionally, incorporating physics-informed nonlinear schedules and exploring alternative statistical criteria represent promising avenues for further advancing generative model efficiency.
+In this paper, we investigated the design of interpolation schedules in flow and diffusion-based generative models within the stochastic interpolants framework.
+We showed that all scalar interpolation schedules are statistically equivalent under the Kullback–Leibler divergence in path space after optimal tuning of diffusion coefficients, yet can exhibit markedly different numerical behavior.
+This equivalence result reveals a fundamental limitation of scalar schedule optimization from a statistical perspective and suggests that breaking this barrier will require more expressive choices, such as matrix-valued, nonlinear, or instance-dependent schedules.
+To leverage the numerical differences among statistically equivalent scalar schedules, we proposed minimizing the averaged squared Lipschitzness of the drift field as a principled criterion for schedule design.
+Unlike kinetic energy minimization in optimal transport, this criterion directly targets numerical regularity and favors dynamics that can be accurately integrated with few time steps.
+Our analytical results show exponential improvements in Lipschitz constants for Gaussian targets and substantial mitigation of mode collapse for Gaussian mixtures.
+Importantly, these insights, derived from analytically tractable settings, transfer effectively to complex high-dimensional invariant measures arising from stochastic Allen–Cahn and Navier–Stokes equations, where they yield robust accuracy improvements across spatial resolutions.
+Overall, our results demonstrate that carefully designed scalar schedules can deliver meaningful practical gains across a wide range of generative modeling tasks, including challenging stochastic PDE settings.
+At the same time, the statistical equivalence of scalar schedules highlights the need to move beyond this class in order to achieve further advances.
+Exploring richer interpolation mechanisms and alternative statistical criteria, as well as understanding their interaction with numerical discretization and physical structure preservation, represents a promising direction for future work in generative modeling for scientific machine learning.
 Acknowledgments YC acknowledges support from UCLA start-up funds.
@@
 Then $I_{t}$ is a Gaussian noisy version of $\beta_{t}x_{1}$, implying that
-$\rho_{t}(x)\propto\int_{\mathbb{R}^{d}}\rho_{t}(y)\exp(-\frac{\|x-y\|_{2}^{2}}{2\alpha_{t}^{2}}){\rm d}y\,.$
+$\rho_{t}(x)\propto\int_{\mathbb{R}^{d}}q_{t}(y)\exp(-\frac{\|x-y\|_{2}^{2}}{2\alpha_{t}^{2}}){\rm d}y\,.$
 Taking gradient yields the formula
-$\nabla\log\rho_{t}(x)=\frac{1}{\int_{\mathbb{R}^{d}}\rho_{t}(y)\exp(-\frac{\|x-y\|_{2}^{2}}{2\alpha_{t}^{2}}){\rm d}y}\int_{\mathbb{R}^{d}}(-\frac{x-y}{\alpha_{t}^{2}})\rho_{t}(y)\exp(-\frac{\|x-y\|_{2}^{2}}{2\alpha_{t}^{2}}){\rm d}y\,.$
+$\nabla\log\rho_{t}(x)=\frac{1}{\int_{\mathbb{R}^{d}}q_{t}(y)\exp(-\frac{\|x-y\|_{2}^{2}}{2\alpha_{t}^{2}}){\rm d}y}\int_{\mathbb{R}^{d}}(-\frac{x-y}{\alpha_{t}^{2}})q_{t}(y)\exp(-\frac{\|x-y\|_{2}^{2}}{2\alpha_{t}^{2}}){\rm d}y\,.$
 On the other hand, by the Bayes rule, we know that
-$\frac{1}{\int_{\mathbb{R}^{d}}\rho_{t}(y)\exp(-\frac{\|x-y\|_{2}^{2}}{2\alpha_{t}^{2}}){\rm d}y}\rho_{t}(y)\exp(-\frac{\|x-y\|_{2}^{2}}{2\alpha_{t}^{2}})$
+$\frac{1}{\int_{\mathbb{R}^{d}}q_{t}(y)\exp(-\frac{\|x-y\|_{2}^{2}}{2\alpha_{t}^{2}}){\rm d}y}q_{t}(y)\exp(-\frac{\|x-y\|_{2}^{2}}{2\alpha_{t}^{2}})$
 is the density of the conditional distribution $\beta_{t}x_{1}|\alpha_{t}z+\beta_{t}x_{1}=x$. Therefore,